### v3

# In this paper, we investigate the mixed norm estimates for the operator TTassociated with a dilated plane curve $(ut,u\gamma(t))$, defined by

Abstract
In this paper, we investigate the mixed norm estimates for the operator TTassociated with a dilated plane curve $(ut,u\gamma(t))$, defined by
$Tf(x,u):=\int_{0}^{1}f(x_{1}-ut,x_{2}-u\gamma(t))\,dt,$
where $x:=(x_{1},x_{2})$ and $\gamma$ is a general plane curve satisfying appropriate smoothness and curvature conditions. More precisely, we establish the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L_{u}^{r}(\mathbb{R}^{2}\times[1,2])$ (space-time) estimates for $T$, whenever $(\frac{1}{p},\frac{1}{q})$ satisfy
$\max\left\{0,\frac{1}{2p}-\frac{1}{2r},\frac{3}{p}-\frac{r+2}{r}\right\}<\frac{1}{q}\leq\frac{1}{p}<\frac{r+1}{2r}$
and
$1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0,$
where $r\in[1,\infty]$ and $\omega:=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$. These results are sharp, except for certain borderline cases. Additionally, we examine the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{u}^{r}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])$ (time-space) estimates for $T$, which are especially almost sharp when $p=2$ or $p\in[1,\frac{3}{2}]\cup[4,\infty]$.
1 Introduction
Let $\gamma\in C^{N}(0,1]$ with $N\in\mathbb{N}$ sufficiently large. We define the operator $T$ along the dilated plane curve $(ut,u\gamma(t))$ as
$Tf(x,u):=\int_{0}^{1}f(x_{1}-ut,x_{2}-u\gamma(t))\,dt,$
which has been extensively studied in works such as Hickman [18], Beltran, Guo, Hickman, and Seeger [2], Ko, Lee, and Oh [24], among others. We also consider the maximal operator
$M_{\gamma}f(x):=\sup_{u\in(0,\infty)}\left|Tf(x,u)\right|.$
Bourgain [6] established foundational estimates for maximal averages over planar convex curves. Beltran, Guo, Hickman, and Seeger [2], as well as Ko, Lee, and Oh [24], demonstrated that the maximal average over a curve in three dimensions is bounded on $L_{x}^{p}(\mathbb{R}^{3})$ if and only if $p>3$. For higher-dimensional settings, we refer to Ko, Lee and Oh [25]. Related studies on maximal functions associated with hypersurfaces can be found in [45, 46, 21] and references therein.
A natural question arises concerning the $L_{x}^{p}(\mathbb{R}^{2})$ to $L_{x}^{q}(\mathbb{R}^{2})$ boundedness of $M_{\gamma}$. It is well-established in various contexts that $M_{\gamma}$ is not $L^{p}\rightarrow L^{q}$ bounded unless $p=q$ (see Liu and Yu [35, Remark 1.3]). However, when the supremum is restricted to $u\in[1,2]$, they derived an almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}(\mathbb{R}^{2})$ estimate for $M_{\gamma}$, i.e., the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])$ estimate for $T$. This phenomenon, known as $L^{p}$-improving, is discussed in greater detail in [52, 40, 1, 32].
In this paper, we aim to extend the estimates for the maximal function to mixed norm estimates for $T$. Our primary objective is to establish $L^{p}_{x}(\mathbb{R}^{2})\rightarrow L^{q}_{x}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])$ (space-time) estimates for $T$, which is related to $r$-variation estimates for $T$ (see [4, 30]). When $q=r$, there is already a substantial body of literature addressing this case. For example, when $\gamma(t):=t^{2}$, the $L_{x}^{2}(\mathbb{R}^{2})\rightarrow L_{x,u}^{6}(\mathbb{R}^{2}\times[1,2])$ boundedness of $T$ was essentially established by Strichartz [51] and later refined by Schlag and Sogge [41]. Additional results on $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x,u}^{q}(\mathbb{R}^{2}\times[1,2])$ boundedness for $T$ can be found in Gressman [14, 15]. More recently, Li, Liu, Lou, and Yu [31] established almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x,u}^{q}(\mathbb{R}^{2}\times[1,2])$ estimates for $T$ along a broader class of curves. In higher dimensions, Hickman [18] obtained nearly sharp $L_{x}^{p}(\mathbb{R}^{n})\rightarrow L_{x,u}^{q}(\mathbb{R}^{n}\times[1,2])$ estimates for $T$ along the moment curve $(t,t^{2},\dots,t^{n})$. He also raised the problem of determining the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$ for which mixed norm estimates for $T$ hold.
The main purpose of this paper is to consider the mixed norm estimates for $T$ in two-dimensional case, but with some general plane curves including the parabola $\gamma(t):=t^{2}$. Our first result is the following.
Theorem 1.1.
Assume $\gamma\in C^{N}(0,1]$ with $N\in\mathbb{N}$ sufficiently large, $\lim_{t\rightarrow 0^{+}}\gamma(t)=0$, and $\gamma$ is monotonic on $(0,1]$. Additionally, suppose $\gamma$ satisfies the following conditions:
  1. (i)
    there exist positive constants $\{C^{(j)}_{1}\}_{j=1}^{2}$ such that $|\frac{t^{j}\gamma^{(j)}(t)}{\gamma(t)}|\geq C^{(j)}_{1}$ for any $t\in(0,1]$;
  2. (ii)
    there exist positive constants $\{C^{(j)}_{2}\}_{j=1}^{N}$ such that $|\frac{t^{j}\gamma^{(j)}(t)}{\gamma(t)}|\leq C^{(j)}_{2}$ for any $t\in(0,1]$.
Then, there exists a positive constant $C$ depending only on $\gamma$, $p$, $q$ and $r$, such that for all $f\in L_{x}^{p}(\mathbb{R}^{2})$,
$\left\|Tf\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}\leq C\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $(\frac{1}{p},\frac{1}{q})$ satisfies
$\max\left\{0,\frac{1}{2p}-\frac{1}{2r},\frac{3}{p}-\frac{r+2}{r}\right\}<\frac{1}{q}\leq\frac{1}{p}<\frac{r+1}{2r}$
and
$1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$
with $r\in[1,\infty]$. Moreover, the range of $(\frac{1}{p},\frac{1}{q})$ is sharp except for some borderlines for all $r\in[1,\infty]$. Here and hereafter, $\omega:=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$.
Remark 1.2.
Here are some examples of curves satisfying the conditions (i) and (ii) of Theorem 1.1. We may add a characteristic function $\chi_{(0,\epsilon_{0}]}(t)$ if necessary, where $\epsilon_{0}$ is small enough.
  1. (1)
    $\gamma_{1}(t):=t^{d}$, where $d\in(0,\infty)$ and $d\neq 1$;
  2. (2)
    $\gamma_{2}(t):=a_{d}t^{d}+a_{d+1}t^{d+1}+\cdots+a_{d+m}t^{d+m}$, where $\geq 2, dd\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ , $a_{d}\neq 0$, i.e., $\gamma_{2}$ is a polynomial of degree at least $d$ with no linear term and constant term;
  3. (3)
    $\gamma_{3}(t):=\sum_{i=1}^{d}\beta_{i}t^{\alpha_{i}}$, where $\alpha_{i}\in(0,\infty)$ for all $i=1,2,\cdots,d$, $\min_{i\in\{1,2,\cdots,d\}}\{\alpha_{i}\}_{i=1}^{d}\neq 1$ and $d\in\mathbb{N}$;
  4. (4)
    $\gamma_{4}(t):=1-\sqrt{1-t^{2}}$, $t\sin t$, $t-\sin t$, $1-\cos t$, or $e^{t}-t-1$, or $t^{d}\ln(1+t)$ with $d\in(0,\infty)$;
  5. (5)
    $\gamma_{5}(t)$ is a smooth function on $[0,1]$ satisfying $\gamma(0)=\gamma^{\prime}(0)=\cdots=\gamma^{(d-1)}(0)=0$ and $\gamma^{(d)}(0)\neq 0$, where $d\geq 2$ and $d\in\mathbb{N}$. From Iosevich [22], we know that $\gamma_{5}$ is finite type $d$ at $t=0$. $\gamma_{2}$ and $\gamma_{4}$ provide some special cases of $\gamma_{5}$.
Remark 1.3.
The range of $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$ is illustrated in Figure 1. This range corresponds to the open convex hull of the points $\{\textrm{O, A, B, C, D}\}$ but with the open line segment $(\textrm{O, A})$. From Figure 1, it is evident that the range of $(\frac{1}{p},\frac{1}{q})$ diminishes as $r$ increases. Specifically, when $r=1$, the range encompasses the entire lower triangle except for the boundary lines $\frac{1}{p}=1$ and $\frac{1}{q}=0$. Conversely, when $r=\infty$, the range coincides with the result obtained by Liu and Yu [35]. Additionally, the range of $(\frac{1}{p},\frac{1}{q})$ is observed to shrink as $\omega$ increases. Notably, the condition $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})>0$ does not impose any restrictions on the range of $(\frac{1}{p},\frac{1}{q})$ when $\omega\leq\frac{4(r-1)}{r+4}$.
Figure 1: The range of $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$.
Another motivation for this paper stems from the study of the maximal function $\mathcal{M}$, defined by averaging over the curve $(t,\gamma(t))$, as follows:
$\mathcal{M}f(x):=\sup_{\epsilon\in(0,\infty)}\frac{1}{2\epsilon}\int_{-\epsilon}^{\epsilon}\left|f(x_{1}-t,x_{2}-\gamma(t))\right|\,\mathrm{d}t.$
This operator, along with related families, has been extensively studied. For the case $\gamma(t):=t^{2}$, Nagel, Riviere, and Wainger [38] demonstrated that $\mathcal{M}$ is bounded on $L_{x}^{p}(\mathbb{R}^{2})$ for all $p\in(1,\infty]$. Stein [48] extended this result to homogeneous curves, while Stein and Wainger [49] investigated smooth curves. For more general families of curves, we refer readers to [50, 8, 9].
Let $\bar{\psi}:\mathbb{R}^{+}\rightarrow\mathbb{R}$ be a smooth function supported on $\{t\in\mathbb{R}:\frac{1}{8}\leq t\leq 8\}$, satisfying $0\leq\bar{\psi}(t)\leq 1$ and $\bar{\psi}(t)=1$ on $\{t\in\mathbb{R}:\frac{1}{4}\leq t\leq 4\}$. Throughout this paper, we denote by $P_{k}$ the operator on $\mathbb{R}^{n}$ with multiplier $\bar{\psi}_{k}(|\xi|):=\bar{\psi}(2^{-k}|\xi|)$, where $k\in\mathbb{N}$. A key estimate required to establish Theorem 1.1 is essentially the following almost sharp space-time estimate for the half-wave propagator:
(1.1) $2^{-\frac{1}{2}k}\left\|e^{iu\sqrt{-\Delta}}P_{k}f\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{-\epsilon k}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $\epsilon$ is a positive constant. It is well known that the solution $g$ of the Cauchy problem for the wave equation in $\mathbb{R}^{n}\times\mathbb{R}$
$\begin{cases}(\partial^{2}_{u}-\Delta)g(x,u)=0,~~(x,u)\in\mathbb{R}^{n}\times\mathbb{R};\\ g(x,0)=f(x),~~\partial_{u}g(x,0)=0,\end{cases}$
can be written in terms of the half-wave propagator $e^{iu\sqrt{-\Delta}}f$ as
$g(x,u)=\frac{1}{2}\left(e^{iu\sqrt{-\Delta}}f(x)+e^{-iu\sqrt{-\Delta}}f(x)\right).$
Let $\dot{H}_{x}^{s}(\mathbb{R}^{n})$ denote the homogeneous Sobolev space. The following space-time estimate for the half-wave propagator is of particular interest in harmonic analysis:
(1.2) $\left\|e^{iu\sqrt{-\Delta}}f\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{n}\times\mathbb{R})}\lesssim\|f\|_{\dot{H}_{x}^{s}(\mathbb{R}^{n})}.$
In the present paper, we focus on this estimate for all $f\in L_{x}^{p}(\mathbb{R}^{2})$ with $p\in[1,\infty]$.
The estimate (1.2) and its local version have been studied by numerous authors. For instance, when $r=\infty$, (1.2) represents a maximal function estimate for the half-wave propagator, which is associated with the pointwise convergence of solutions to their initial data in $L_{x}^{2}(\mathbb{R}^{n})$. Cowling [11] proved that (1.2) holds for all $q=2$ and $s>\frac{1}{2}$. Rogers and Villarroya [39] later extended this result to all $q\in(2,\infty]$. For endpoint estimates and historical remarks, see Cho, Lee, and Li [10]. When the initial data lies in Sobolev spaces $W_{x}^{s,p}(\mathbb{R}^{n})$, and $q=p$, (1.2) holds for $p\geq\frac{2(n+1)}{n-1}$ and $s>(n-1)\left|\frac{1}{2}-\frac{1}{p}\right|$, as a corollary of Bourgain and Demeter [7]. The two-dimensional case of this estimate can be found in Guth, Wang, and Zhang [16], while an endpoint result in dimensions $n\geq 4$ was established by Heo, Nazarov, and Seeger [17]. When $r=2$ and $q=p>\frac{2(n+2)}{n}$, (1.2) is closely related to Stein’s square function estimates for Bochner-Riesz operators, as discussed in Lee, Rogers, and Seeger [27]. Additionally, analogous results for Fourier integral operators can be found in [44, 36, 37, 12], and for compact manifolds in Beltran, Hickman, and Sogge [3].
It is also natural to consider the time-space estimate for $T$, commonly referred to as the Strichartz estimate. For the half-wave propagator, this estimate can be expressed as
(1.3) $\left\|e^{iu\sqrt{-\Delta}}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{n}\times\mathbb{R})}\lesssim\|f\|_{\dot{H}_{x}^{s}(\mathbb{R}^{n})},$
where it is well-known that (1.3) holds if $q,r\geq 2$, $\frac{1}{r}+\frac{n}{q}=\frac{n}{2}-s$, and $\frac{2}{r}+\frac{n-1}{q}\leq\frac{n-1}{2}$, with the exception of $(r,q)=(2,\infty)$ when $n=3$. For further details and historical remarks, see Keel and Tao [23]. When $q=\infty$, (1.3) is also connected to a Kakeya-type problem for circles, particularly when the initial data belongs to $W_{x}^{\frac{1}{2},3}(\mathbb{R}^{2})$; see Wolff [53]. Over the past few decades, extensive literature has been devoted to this problem, including [51, 28, 29, 5].
We now state our second result concerning the time-space estimate for $T$.
Theorem 1.4.
Let $\gamma$ be as defined in Theorem 1.1. Then, there exists a positive constant $C$ depending only on $\gamma$, $p$, $q$ and $r$, such that for all $f\in L_{x}^{p}(\mathbb{R}^{2})$,
(1.4) $\left\|Tf\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\leq C\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
provided that $(\frac{1}{p},\frac{1}{q})$ satisfies $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})>0$ and one of the following conditions holds:
  1. $A:$
    For $1\leq p\leq q\leq r\leq\infty$, $(\frac{1}{p},\frac{1}{q})$ satisfies $\tilde{s}_{A,0}(p,q,r)<0$;
  2. $B:$
    For $1\leq p\leq r\leq q\leq\infty$, $(\frac{1}{p},\frac{1}{q})$ satisfies $\tilde{s}_{B,0}(p,q,r)<0$, where
    $\tilde{s}_{B,0}(p,q,r):=\min\left\{\tilde{s}_{A,0}(p,r,r)+2\left(\frac{1}{r}-\frac{1}{q}\right),\tilde{s}_{A,0}(p,q,q)\right\};$
  3. $C:$
    For $1\leq r<p\leq q\leq\infty$, $(\frac{1}{p},\frac{1}{q})$ satisfies $\tilde{s}_{C,0}(p,q,r)<0$, where
    $\tilde{s}_{C,0}(p,q,r):=\min\left\{\tilde{s}_{A,0}(p,p,p)+2\left(\frac{1}{p}-\frac{1}{q}\right),\tilde{s}_{A,0}(p,q,q)\right\}.$
Conversely, (1.4) holds only if $(\frac{1}{p},\frac{1}{q})$ satisfies the following conditions:
  1. (I)
    $\frac{1}{q}\leq\frac{1}{p}$;
  2. (II)
    $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})\geq 0$;
  3. (III)
    $\frac{1}{q}\geq\max\left\{\frac{2}{p}-1,\frac{1}{p}-\frac{1}{3}\right\}$;
  4. (IV)
    $\frac{1}{q}\geq\max\left\{\frac{1}{2p}-\frac{1}{2r},\frac{2}{p}-\frac{r+1}{r}\right\}$ and $\max\left\{\frac{1}{r}-\frac{1}{q},0\right\}\geq\tilde{s}_{A,0}(p,q,q)$ for all $r\in[1,\infty]$.
Here, for any $1\leq p,q,r\leq\infty$, we define
$\tilde{s}_{A,0}(p,q,r):=\begin{cases}\frac{1}{p}-\frac{2}{q}-\frac{1}{r},&\text{for }q\geq 3p^{\prime};\\ \frac{1}{p}-\frac{2}{q}+\left(\frac{1}{2p}+\frac{1}{2q}-\frac{1}{2}\right)\frac{q}{r},&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-\frac{1}{q}-1,&\text{for }q\leq p^{\prime}.\end{cases}$
Remark 1.5.
From the definition of $\tilde{s}_{A,0}$ with $p^{\prime}<q<3p^{\prime}$, it is evident that Theorem 1.4 is not always sharp, since the indices $\frac{1}{p}$, $\frac{1}{q}$ and $\frac{1}{r}$ in $\tilde{s}_{A,0}$ is not a linear combination. However, by interpolating between the vertices of the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$, we may obtain some sharp results. It is natural to investigate what occurs in Theorem 1.4 when $p=2$. By interpolating between the points $(\frac{1}{q},\frac{1}{r})=(\frac{1}{6},\frac{1}{6})$ and $(\frac{1}{4},0)$, we obtain that
(1.5) $\left\|Tf\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}$
holds if $(\frac{1}{q},\frac{1}{r})$ lies on the closed line segment $(\frac{1}{2},1),(\frac{1}{2},0)]$ or within the open pentagon with vertices $(\frac{1}{6},\frac{1}{6})$, $(\frac{1}{2},1)$, $(\frac{1}{2},0)$, and $(\frac{1}{4},0)$. Moreover, this estimate is sharp except the boundary lines. The range of $(\frac{1}{q},\frac{1}{r})$ is illustrated in Figure 2.
Figure 2: The range of $(\frac{1}{q},\frac{1}{r})$ in Theorem 1.4 when $p=2$.
Furthermore, a direct computation reveals that Theorem 1.4 is almost sharp for all $p\in[1,\frac{3}{2}]\cup[4,\infty]$. The corresponding range of $(\frac{1}{q},\frac{1}{r})$ can be found in Figure 3.
Figure 3: The range of $(\frac{1}{q},\frac{1}{r})$ in Theorem 1.4 for $p\in[1,\frac{3}{2}]$ and $p\in[4,\infty]$.
Remark 1.6.
We present the detailed result in Theorem 1.4 with $r=2$. Combining Theorem 1.4 with Remark 4.1, if $\left(\frac{1}{p},\frac{1}{q}\right)$ lies on the open line segment $((0,0),(1,1))$ or within the open (dashed) pentagon with vertices $(0,0)$, $(1,1)$, $\left(\frac{2}{3},\frac{1}{3}\right)$, $\left(\frac{1}{2},\frac{1}{6}\right)$, and $\left(\frac{1}{4},0\right)$, we have
(1.6) $\left\|Tf\right\|_{L^{2}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Conversely, (1.6) holds only if $\left(\frac{1}{p},\frac{1}{q}\right)$ belongs to the closed (black) quadrilateral with vertices $(0,0)$, $(1,1)$, $\left(\frac{2}{3},\frac{1}{3}\right)$, and $\left(\frac{1}{3},0\right)$, see Figure 4. We note that if $\gamma(t):=t^{d}$ with $d\geq 3$, the time-space estimate $T:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{2}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ established in Theorem 1.4 is essentially sharp. However, if $\gamma(t):=t^{d}$ with $d\in(0,1)\cup(1,2)$, the condition $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$ is redundant.
Figure 4: The range of $\left(\frac{1}{p},\frac{1}{q}\right)$ in Theorem 1.4 with $r=2$.
The main contribution of this paper is to establish the optimal range of $\left(\frac{1}{p},\frac{1}{q},\frac{1}{r}\right)$ for the space-time estimate of $T$, as stated in Theorem 1.1. Indeed, the range of $\left(\frac{1}{p},\frac{1}{q}\right)$ for $T:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])$ determines all other ranges of $\left(\frac{1}{p},\frac{1}{q},\frac{1}{r}\right)$. To achieve this result we need the sharp estimate for
$F_{j,k}f(x,u):=\Omega(x,u)\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}e^{i\varphi_{j}(u,\xi,t_{0})}m_{j,k}(u,\xi)\hat{f}(\xi)\,d\xi,$
defined in (2.9), which is essentially a Fourier integral operator with phase function $x\cdot\xi-u|\xi|$. Thus it is basically a half-wave propagator. These estimates for $F_{j,k}$ generalize the results of Mockenhaupt, Seeger, and Sogge [36, Theorem 3.1], and Lee [26, Corollary 1.5] when we restrict our attention to $P_{k}$.
It is evident that the time-space estimate for $T$ differs from the corresponding space-time estimate. Firstly, the latter is translation-invariant when the temporal integral is evaluated before the spatial integral, but not the former. Secondly, as stated in Subsection 4.1, the time-space estimate for $T$ not always can be reduced to the corresponding time-space estimate similar to (1.1). Thirdly, we can only obtain a common estimate
$\|F_{j,k}f\|_{L^{r}_{u}L_{x}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{k}\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}$
for all $r\in[1,\infty]$, not as the corresponding space-time estimate can be improved directly when $r=1$; see (3.3). Our main tools will be interpolation, a fixed time estimate for $F_{j,k}$ based on Seeger, Sogge and Stein [42, Theorem 2.1], Gao, Liu, Miao and Xi [12, Theorem 1.4] and some results stated in Stein [47, Chapters VIII and IX].
To establish the necessity part in Theorem 1.4, we utilized the sharp range of $\left(\frac{1}{p},\frac{1}{q}\right)$ in $T:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x,u}^{q}(\mathbb{R}^{2}\times[1,2])$, not only the sharp range of $(\frac{1}{p},\frac{1}{q})$ in $T:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])$. We also constructed some examples inspired by previous papers. However, it should be noted that Theorem 1.4 may not be sharp in some case of $p\in\left(\frac{3}{2},2\right)\cup(2,4)$. To be more precise, it follows from Theorem 1.4 that
(1.7) $\left\|Tf\right\|_{L^{p}_{u}L_{x}^{\infty}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
holds if $\max\{1+\omega,4\}<p\leq\infty$ and only if $\max\{1+\omega,3\}\leq p\leq\infty$. Thus the range of $(\frac{1}{p},\frac{1}{q},\frac{1}{r})$ in (1.4) is not sharp. Moreover, we can conjecture that the infimum of $p$ in (1.7) is $\max\{1+\omega,3\}$ by the estimate $(2)$ established in Wolff [53]. From Theorem 1.4, we also note that (1.7) is essentially sharp if $\gamma(t):=t^{d}$ with $d\geq 3$.
This paper is organized as follows: In Section 2, we provide preliminaries based on prior papers [35, 31] and reduce our estimates to a Fourier integral operator $F_{j,k}$. In Section 3, we prove Theorem 1.1 by addressing the sufficiency and necessity in Subsections 3.1 and 3.2, respectively. In Section 4, we discuss the time-space estimate for $T$, i.e., Theorem 1.4. Specifically, Subsection 4.1 elaborates on the estimate (1.4) by analyzing three cases: $1\leq p\leq q\leq r\leq\infty$, $1\leq p\leq r\leq q\leq\infty$, and $1\leq r<p\leq q\leq\infty$. The necessary conditions are verified in Subsection 4.2.
Finally, we establish some notation conventions. Throughout this paper, the letter $C$ denotes a positive constant, independent of the essential variables, but whose value may vary from line to line. We use $C_{(u,N,\alpha,\ldots)}$ to denote a positive constant depending on the parameters $u,N,\alpha,\ldots$. The notation $a\lesssim b$ (or $a\gtrsim b$) indicates that there exists a finite positive constant $C$ such that $a\leq Cb$ (or $a\geq Cb$). $a\approx b$ means both $a\lesssim b$ and $b\lesssim a$ hold. For $x\in\mathbb{R}^{n}$ and $r\in(0,\infty)$, $B(x,r):=\{y\in\mathbb{R}^{n}:|x-y|<r\}$ and $B^{\complement}(x,r)$ is its complement in $\mathbb{R}^{n}$. Let $\mathbb{S}^{n-1}$ denote the unit sphere on $\mathbb{R}^{n}$. $\mathcal{S}(\mathbb{R}^{n})$ denotes the collection of all Schwartz functions on $\mathbb{R}^{n}$. $\hat{f}$ and $f^{\vee}$ denote the Fourier transform and inverse Fourier transform of $f$, respectively. For $1<q\leq\infty$, $q^{\prime}$ denotes the adjoint number of $q$, i.e., $\frac{1}{q}+\frac{1}{q^{\prime}}=1$. $\mathbb{Z}_{0}^{-}:=\{0,-1,-2,\ldots\}$, $\mathbb{N}:=\{1,2,\ldots\}$, $\mathbb{N}_{0}:=\{0,1,2,\ldots\}$, and $\mathbb{R}^{+}:=(0,\infty)$. For any set $E$, $\chi_{E}$ denotes the characteristic function of $E$.
2 Preliminaries
The proofs of Theorems 1.1 and 1.4 rely on some results established in previous papers [35, 31]. For ease of reading, we provide a sketch of reducing the estimate to a Fourier integral operator in this section. Detailed discussions can be found in [35, 31]. We begin by collecting some lemmas.
Lemma 2.1.
([33, Lemma 2.2] or [34, Lemma 2.2]) Let $\gamma$ be defined as in Theorem 1.1 and let $\Gamma_{j}(t):=\frac{\gamma(2^{j}t)}{\gamma(2^{j})}$ with $j\in\mathbb{Z}_{0}^{-}$ and $t\in[\frac{1}{2},2]$. The following inequalities hold uniformly in $j$:
  1. (i)
    $e^{-C^{(1)}_{2}}\leq\Gamma_{j}(t)\leq e^{C^{(1)}_{2}}$;
  2. (ii)
    $\frac{C^{(1)}_{1}}{2e^{C^{(1)}_{2}}}\leq|\Gamma_{j}^{\prime}(t)|\leq 2e^{C^{(1)}_{2}}C^{(1)}_{2}$;
  3. (iii)
    $\frac{C^{(2)}_{1}}{4e^{C^{(1)}_{2}}}\leq|\Gamma_{j}^{\prime\prime}(t)|\leq 4e^{C^{(1)}_{2}}C^{(2)}_{2}$;
  4. (iv)
    $|\Gamma_{j}^{(k)}(t)|\leq 2^{k}e^{C^{(1)}_{2}}C^{(k)}_{2}$ for all $2\leq k\leq N$ and $k\in\mathbb{N}$;
  5. (v)
    $|((\Gamma_{j}^{\prime})^{-1})^{(k)}(t)|\lesssim 1$ for all $0\leq k<N$ and $k\in\mathbb{N}$, where $(\Gamma_{j}^{\prime})^{-1}$ is the inverse function of $\Gamma_{j}^{\prime}$.
Lemma 2.2.
([43, Lemma 2.4.2]) Suppose that $F$ is $C^{1}(\mathbb{R})$. Then, if $q>1$,
$\sup_{u\in[1,2]}|F(u)|^{q}\leq|F(1)|^{q}+q\left(\int_{1}^{2}|F(u)|^{q}\,\mathrm{d}u\right)^{\frac{1}{q^{\prime}}}\left(\int_{1}^{2}|F^{\prime}(u)|^{q}\,\mathrm{d}u\right)^{\frac{1}{q}}.$
Lemma 2.3.
([35, Lemma 2.3]) Recall that $\omega=\limsup_{t\rightarrow 0^{+}}\frac{\ln|\gamma(t)|}{\ln t}$. Then, for all $(\frac{1}{p},\frac{1}{q})$ satisfying $\frac{1}{q}\leq\frac{1}{p}$ and $1+(1+\omega)(\frac{1}{q}-\frac{1}{p})>0$, we have
$\sum_{j\in\mathbb{Z}_{0}^{-}}2^{j}|2^{j}\gamma(2^{j})|^{\frac{1}{q}-\frac{1}{p}}<\infty.$
We consider the space-time estimate for the operator
(2.1) $Tf(x,u)=\int_{0}^{1}f(x_{1}-ut,x_{2}-u\gamma(t))\,\mathrm{d}t,$
where, without loss of generality, we assume $f\geq 0$. Let $\psi$ be a smooth and positive function supported on the interval $\{t\in\mathbb{R}:\frac{1}{2}\leq t\leq 2\}$ such that $\sum_{j\in\mathbb{Z}}\psi_{j}(t)=1$ for all $t>0$, where $\psi_{j}(t):=\psi(2^{-j}t)$. For any $j\in\mathbb{Z}$, we define $\delta_{j}f(x_{1},x_{2}):=f(2^{j}x_{1},\gamma(2^{j})x_{2})$. Note that $|2^{j}\gamma(2^{j})|^{\frac{1}{q}}\|\delta_{j}f\|_{L^{q}(\mathbb{R}^{2})}=\|f\|_{L^{q}(\mathbb{R}^{2})}$, it suffices to show that
$\sum_{j\in\mathbb{Z}_{0}^{-}}2^{j}|2^{j}\gamma(2^{j})|^{\frac{1}{q}-\frac{1}{p}}\left\|\widetilde{T}_{j}\right\|_{L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 1,$
where
$\widetilde{T}_{j}f(x,u):=\int_{0}^{\infty}f(x_{1}-ut,x_{2}-u\Gamma_{j}(t))\psi(t)\,\mathrm{d}t,$
and the definition of $\Gamma_{j}$ can be found in Lemma 2.1. Consequently, by Lemma 2.3, the proof reduces to establishing the estimate $\widetilde{T}_{j}:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])$ uniformly in $j\in\mathbb{Z}_{0}^{-}$.
Let $\Psi^{0}:=\sum_{k\leq 0}\psi_{k}$ and define
$H_{j}(u,\xi):=\int_{\mathbb{R}}e^{-iu\xi_{1}t-iu\xi_{2}\Gamma_{j}(t)}\psi(t)\,\mathrm{d}t,$
where $\xi:=(\xi_{1},\xi_{2})$. We decompose $\widetilde{T}_{j}$ as
$\widetilde{T}_{j}f(x,u)=\widetilde{T}^{0}_{j}f(x,u)+\sum_{k\geq 1}\widetilde{T}_{j,k}f(x,u),$
where the multipliers of $\widetilde{T}^{0}_{j}$ and $\widetilde{T}_{j,k}$ are $H_{j}(u,\xi)\Psi^{0}(u|\xi|)$ and $H_{j}(u,\xi)\psi_{k}(u|\xi|)$, respectively. For $\widetilde{T}^{0}_{j}$, by Lemma 2.1, we can obtain that $|\partial_{u}^{\alpha}\partial^{\beta}_{\xi}(H_{j}(u,\xi)\Psi^{0}(u|\xi|))|\lesssim(1+|\xi|)^{-3}$ for all $(\alpha,\beta)\in\mathbb{N}_{0}\times\mathbb{N}^{2}_{0}$ with $|\alpha|\leq 1$ and $|\beta|\leq 3$, where the implicit constant is independent of $u\in[1,2]$ and $j\in\mathbb{Z}_{0}^{-}$. Furthermore, by Lemma 2.2, Hölder’s inequality, and Young’s inequality, it follows that
(2.2) $\left\|\widetilde{T}^{0}_{j}f\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}\leq\left\|\widetilde{T}^{0}_{j}f\right\|_{L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
for all $1\leq r\leq\infty$ and $1\leq p\leq q\leq\infty$.
For $\widetilde{T}_{j,k}$, we further split it into $\widetilde{T}^{1}_{j,k}$ and $\widetilde{T}^{2}_{j,k}$, where the multipliers of $\widetilde{T}^{1}_{j,k}$ and $\widetilde{T}^{2}_{j,k}$ are
$\rho\left(\frac{|\xi_{1}|}{|\xi_{2}|}\right)H_{j}(u,\xi)\psi_{k}(u|\xi|)~~\textrm{and}~~\left(1-\rho\left(\frac{|\xi_{1}|}{|\xi_{2}|}\right)\right)H_{j}(u,\xi)\psi_{k}(u|\xi|),$
respectively. Here, $\rho\in C^{\infty}_{c}(\mathbb{R}^{+})$ is a smooth cutoff function satisfying $\rho=1$ on the interval $\frac{C^{(1)}_{1}}{10e^{C^{(1)}_{2}}},10e^{C^{(1)}_{2}}C^{(1)}_{2}$]. Then, as in (2.2), by Lemmas 2.1 and 2.2, we obtain the decay estimate
$\left\|\widetilde{T}^{2}_{j,k}f\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{-\frac{k}{q^{\prime}}}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
for all $1\leq r\leq\infty$ and $1\leq p\leq q\leq\infty$. Therefore, it suffices to analyze the operator $\widetilde{T}^{1}_{j,k}$.
Let the phase function in $H_{j}(u,\xi)$ be $\varphi_{j}(u,\xi,t):=-u\xi_{1}t-u\xi_{2}\Gamma_{j}(t)$. We focus on the case where there exists a critical point $t_{0}\in\left(\frac{1}{2},2\right)$ of $\varphi_{j}$. By the method of stationary phase, we write
(2.3) $\widetilde{T}^{1}_{j,k}f(x,u)=\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}e^{i\varphi_{j}(u,\xi,t_{0})}m_{j,k}(u,\xi)\hat{f}(\xi)\,\mathrm{d}\xi,$
where
$m_{j,k}(u,\xi):=\rho\left(\frac{|\xi_{1}|}{|\xi_{2}|}\right)\psi_{k}(u|\xi|)\int_{\mathbb{R}}e^{-iu\xi_{2}t^{2}\eta_{j}(t,t_{0})}\psi(t+t_{0})\,\mathrm{d}t,$
with $\eta_{j}(t,t_{0}):=\frac{\Gamma^{\prime\prime}_{j}(t_{0})}{2}+\frac{t}{2}\int_{0}^{1}(1-\theta)^{2}\Gamma^{\prime\prime\prime}_{j}(\theta t+t_{0})\,\mathrm{d}\theta$. Moreover, for all $(\alpha,\beta)\in\mathbb{N}^{2}_{0}\times\mathbb{N}_{0}$ with $|\alpha|+|\beta|<N$,
(2.4) $\left|\partial_{u}^{\beta}\partial_{\xi}^{\alpha}m_{j,k}(u,\xi)\right|\lesssim(1+|\xi|)^{-\frac{1}{2}-|\alpha|}$
uniformly in $u\in(\frac{1}{2},\frac{5}{2})$.
We can also express $\widetilde{T}^{1}_{j,k}$ as
$\widetilde{T}^{1}_{j,k}f(x,u)=\int_{\mathbb{R}^{2}}\widetilde{K}_{j,k}(x-y,u)f(y)\,\mathrm{d}y,$
where the kernel is given by
(2.5) $\widetilde{K}_{j,k}(x,u):=\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}e^{i\varphi_{j}(u,\xi,t_{0})}m_{j,k}(u,\xi)\,\mathrm{d}\xi.$
Furthermore, we split $\widetilde{T}^{1}_{j,k}$ into the sum of $\widetilde{T}^{1,a}_{j,k}$ and $\widetilde{T}^{1,b}_{j,k}$, whose kernels are $\chi_{B^{\complement}(0,\varpi)}(x)\widetilde{K}_{j,k}(x,u)$ and $\chi_{B(0,\varpi)}(x)\widetilde{K}_{j,k}(x,u),$ respectively. Here, $\varpi$ is a positive constant chosen large enough such that
$\left|\nabla_{\xi}[(x-y)\cdot\xi+\varphi_{j}(u,\xi,t_{0})]\right|\gtrsim|x-y|$
when $|x-y|\geq\varpi$. The desired decay estimate for $\widetilde{T}^{1,a}_{j,k}$ can be obtained via integration by parts. It then suffices to consider $\widetilde{T}^{1,b}_{j,k}$.
Let $\Omega:\mathbb{R}^{2}\times\mathbb{R}\rightarrow\mathbb{R}$ be a nonnegative smooth function such that $\Omega(x,u)=1$ on
$\{(x,u)\in\mathbb{R}^{2}\times\mathbb{R}:|x|\leq\varpi,~u\in[1,2]\}$
and vanishes outside
$\{(x,u)\in\mathbb{R}^{2}\times\mathbb{R}:|x|\leq 2\varpi,~u\in(\tfrac{1}{2},\tfrac{5}{2})\}.$
Define
(2.6) $F_{j,k}f(x,u):=\Omega(x,u)\int_{\mathbb{R}^{2}}\widetilde{K}_{j,k}(x-y,u)f(y)\,\mathrm{d}y$
and
(2.7) $E_{j,k}f(x,u):=\Omega(x,u)\int_{\mathbb{R}^{2}}\chi_{B(0,\varpi)}(x-y)\widetilde{K}_{j,k}(x-y,u)f(y)\,\mathrm{d}y.$
To estimate $\widetilde{T}^{1,b}_{j,k}$, as in [35], it suffices to obtain a corresponding decay space-time estimate for $F_{j,k}$. Indeed, we partition $\mathbb{R}^{2}$ into $\bigcup_{i\in\mathbb{Z}^{2}}B(x_{i},\varpi)$ such that $|x_{i}-x_{i^{\prime}}|\approx|i-i^{\prime}|\varpi$ for any $i\neq i^{\prime}$. Then, $\|\widetilde{T}^{1,b}_{j,k}f\|^{q}_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}$ is bounded from above by
(2.8) $\sum_{i\in\mathbb{Z}^{2}}\left\|\left(E_{j,k}-F_{j,k}\right)\left(f(\cdot-x_{i})\right)\right\|^{q}_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times\mathbb{R})}+\sum_{i\in\mathbb{Z}^{2}}\left\|F_{j,k}\left(f(\cdot-x_{i})\right)\right\|^{q}_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times\mathbb{R})}.$
The first term can be handled by a similar argument as for $\widetilde{T}^{1,a}_{j,k}$, and the second term reduces to a corresponding decay space-time estimate for $F_{j,k}$.
Putting things together, we have reduced our estimate to $F_{j,k}$. We note that
(2.9) $F_{j,k}f(x,u):=\Omega(x,u)\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}e^{i\varphi_{j}(u,\xi,t_{0})}m_{j,k}(u,\xi)\hat{f}(\xi)\,\mathrm{d}\xi,$
which is a Fourier integral operator of order $\frac{1}{2}$ and satisfies the special assumption that
(2.10) $\text{rank}~\partial^{2}_{\xi\xi}\left(x\cdot\xi+\varphi_{j}(u,\xi,t_{0})\right)=1$
when $\xi\neq 0$.
3 Proof of Theorem 1.1
In this section, we present the proof of Theorem 1.1, building upon the preliminary results established in Section 2. The proof is divided into two parts.
3.1 Sufficiency of the region $(\frac{1}{p},\frac{1}{q})$ for all $r\in[1,\infty]$
We begin the proof with the following two lemmas.
Lemma 3.1.
For $\widetilde{K}_{j,k}$ defined in (2.5) and $N\in\mathbb{N}$ sufficiently large, for all $\alpha\in\mathbb{N}_{0}$, there exists a positive constant $C_{(N,\alpha)}$, depending on $N$ and $\alpha$, such that
$\left|\partial_{u}^{\alpha}\widetilde{K}_{j,k}(x,u)\right|\leq C_{(N,\alpha)}\frac{2^{(1+\alpha)k}}{\left(1+2^{k}\left||x|-u\right|\right)^{N}}$
holds for all $x\in\mathbb{R}^{2}$ and $u\in(\frac{1}{2},\frac{5}{2})$.
Proof.
Note that $t_{0}=(\Gamma^{\prime}_{j})^{-1}(-\frac{\xi_{1}}{\xi_{2}})\in(\frac{1}{2},2)$. By a change of variables, we can express
$\widetilde{K}_{j,k}(x,u)=2^{2k}\int_{\mathbb{R}^{2}}e^{i2^{k}x\cdot\xi}e^{i2^{k}\varphi_{j}(u,\xi,t_{0})}m_{j,k}(u,2^{k}\xi)\,\mathrm{d}\xi.$
From $\partial_{\xi_{1}}t_{0}=-\frac{1}{\xi_{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})}$ and $\partial_{\xi_{2}}t_{0}=\frac{\xi_{1}}{\xi_{2}^{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})}$, a straightforward calculation yields
$\left\{\begin{aligned} &\partial_{\xi_{1}\xi_{1}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=u\frac{1}{\xi_{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})};\\ &\partial_{\xi_{1}\xi_{2}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=-u\frac{\xi_{1}}{\xi_{2}^{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})};\\ &\partial_{\xi_{2}\xi_{1}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=-u\frac{\xi_{1}}{\xi_{2}^{2}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})};\\ &\partial_{\xi_{2}\xi_{2}}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=u\frac{\xi_{1}^{2}}{\xi_{2}^{3}}\frac{1}{\Gamma^{\prime\prime}_{j}(t_{0})}.\end{aligned}\right$.
Combining this with Lemma 2.1 and the facts that $|\xi|\approx 1$ and $|\xi_{1}|\approx|\xi_{2}|$, we obtain
$\left|\partial^{\beta}_{\xi}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))\right|\gtrsim 1$
for all multi-indices $\beta$ with $|\beta|=2$. Noting that $\partial_{\xi}(x\cdot\xi+\varphi_{j}(u,\xi,t_{0}))=x-u(t_{0},\Gamma_{j}(t_{0}))$ and $|(t_{0},\Gamma_{j}(t_{0}))|\approx 1$, integration by parts (using Stein [47, Chapter VIII, Section 2, Proposition 5] and (2.4)) yields
$\left|\widetilde{K}_{j,k}(x,u)\right|\lesssim\frac{2^{k}}{\left(1+2^{k}\left||x|-u\right|\right)^{N}}.$
A simple calculation provides the corresponding estimate for $|\partial_{u}^{\alpha}\widetilde{K}_{j,k}(x,u)|$. This completes the proof of the lemma. ∎
Lemma 3.2.
Suppose $G(x,u)\in C_{c}^{1}(\frac{1}{2},\frac{5}{2})$ with respect to $u$ for all $x\in\mathbb{R}^{2}$. Then, for any $1\leq q,r\leq\infty$,
$\left\|G\right\|_{L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|G\right\|^{\frac{1}{r^{\prime}}}_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\left\|\partial_{u}G\right\|^{\frac{1}{r}}_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times\mathbb{R})}.$
Proof.
For $r=\infty$, the result is trivial. For $r\in[1,\infty)$, note that
$|G(x,u)|^{r}\lesssim\int_{\mathbb{R}}|G(x,v)|^{r-1}|\partial_{v}G(x,v)|\,\mathrm{d}v$
holds for all $u\in(\frac{1}{2},\frac{5}{2})$ and $x\in\mathbb{R}^{2}$. By Hölder’s inequality,
$\left\|G(x,u)\right\|_{L^{\infty}_{u}(\mathbb{R})}\lesssim\left(\left\|G(x,u)\right\|^{r-1}_{L^{r}_{u}(\mathbb{R})}\left\|\partial_{u}G(x,u)\right\|_{L^{r}_{u}(\mathbb{R})}\right)^{\frac{1}{r}}$
for all $x\in\mathbb{R}^{2}$. Applying Hölder’s inequality again yields the desired estimate. This completes the proof. ∎
We now proceed to prove the sufficiency part of Theorem 1.1. From Lemma 3.1, it follows that $\|\widetilde{K}_{j,k}\|_{L_{x}^{1}(\mathbb{R}^{2})}\lesssim 1$ uniformly in $u\in(\frac{1}{2},\frac{5}{2})$. Combining this with Young’s inequality, we obtain
(3.1) $\left\|F_{j,k}f\right\|_{L_{x}^{\infty}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{\infty}(\mathbb{R}^{2})}.$
Additionally, Lemma 3.1 implies
(3.2) $\|F_{j,k}f\|_{L_{x}^{\infty}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{k}\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}.$
This bound can be refined by considering the $L^{1}$-norm in $u$. Specifically, we observe that
$\int_{1}^{2}\frac{2^{k}}{\left(1+2^{k}\left||x-y|-u\right|\right)^{N}}\,\mathrm{d}u\lesssim 1$
uniformly in $x,y\in\mathbb{R}^{2}$, which further implies
(3.3) $\left\|F_{j,k}f\right\|_{L_{x}^{\infty}L_{u}^{1}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}.$
Similarly, by
$\int_{\mathbb{R}^{2}}\frac{2^{k}}{\left(1+2^{k}\left||x-y|-u\right|\right)^{N}}\,\mathrm{d}x\lesssim 1$
uniformly in $y\in\mathbb{R}^{2}$ and $u\in(\frac{1}{2},\frac{5}{2})$, we conclude that
(3.4) $\left\|F_{j,k}f\right\|_{L_{x}^{1}L_{u}^{1}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}.$
From [35, (2.44)], for any $\epsilon\in(0,\infty)$, we have
$\left\|F_{j,k}f\right\|_{L_{x}^{5}L_{u}^{5}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{(\epsilon-\frac{1}{5})k}\|f\|_{L_{x}^{\frac{5}{2}}(\mathbb{R}^{2})}.$
Furthermore, by Lemma 2.2 or Lemma 3.2, we find that
(3.5) $\left\|F_{j,k}f\right\|_{L_{x}^{5}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\epsilon k}\|f\|_{L_{x}^{\frac{5}{2}}(\mathbb{R}^{2})}$
holds for any $\epsilon\in(0,\infty)$. On the other hand, by (2.4), it is a consequence of Plancherel’s theorem that
(3.6) $\left\|F_{j,k}f\right\|_{L_{x}^{2}L_{u}^{2}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{-\frac{1}{2}k}\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}.$
Combining this with Lemma 2.2 and (2.4), we obtain
(3.7) $\begin{split}\left\|F_{j,k}f\right\|_{L_{x}^{2}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}&\lesssim\left\|F_{j,k}f(x,1)\right\|_{L_{x}^{2}(\mathbb{R}^{2})}+\left(\left\|F_{j,k}f\right\|_{L_{x}^{2}L_{u}^{2}(\mathbb{R}^{2}\times\mathbb{R})}\left\|\partial_{u}(F_{j,k}f)\right\|_{L_{x}^{2}L_{u}^{2}(\mathbb{R}^{2}\times\mathbb{R})}\right)^{\frac{1}{2}}\\ &\lesssim\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}.\end{split}$
From Figure 1, for all $r\in[1,\infty]$, it suffices to obtain a constant bound $C$ or a small bound $2^{\epsilon_{1}k}$ at points $\{\text{O, A, B, C, D}\}$ for any $\epsilon_{1}\in(0,\infty)$, and a decay bound $2^{-\epsilon_{2}k}$ for $F_{j,k}$ at $(\frac{1}{4},\frac{1}{4})$ with some $\epsilon_{2}\in(0,\infty)$. More precisely, we first claim the following:
  1. $\bullet$
    Constant bound at point O: For any $r\in[1,\infty]$, by Hölder’s inequality and (3.1), it is easy to see that
    $\left\|F_{j,k}f\right\|_{L_{x}^{\infty}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{\infty}(\mathbb{R}^{2})}.$
  2. $\bullet$
    Constant bound at point A: It follows from interpolation between the $L_{x}^{1}(\mathbb{R}^{2})\rightarrow L_{x}^{1}L_{u}^{1}(\mathbb{R}^{2}\times\mathbb{R})$ and $L_{x}^{2}(\mathbb{R}^{2})\rightarrow L_{x}^{2}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})$ estimates of $F_{j,k}$ established in (3.4) and (3.7), respectively.
  3. $\bullet$
    Constant bound at point B: For any $r\in[1,\infty]$, from (3.7) and (3.3), reiterating this interpolation process, then the following estimate is valid:
    $\left\|F_{j,k}f\right\|_{L_{x}^{\frac{2r}{r-1}}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{\frac{2r}{r+1}}(\mathbb{R}^{2})}.$
    This is the desired estimate at the point B.
  4. $\bullet$
    Small bound at point C: For any $r\in[1,\infty]$, a further interpolation between (3.5) and (3.3) gives
    $\left\|F_{j,k}f\right\|_{L_{x}^{\frac{5r}{r-1}}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\frac{\epsilon}{r^{\prime}}k}\|f\|_{L_{x}^{\frac{5r}{2r+3}}(\mathbb{R}^{2})}$
    for any $\epsilon\in(0,\infty)$.
  5. $\bullet$
    Constant bound at point D: For any $r\in[1,\infty]$, interpolation between (3.1) and (3.3) yields
    $\left\|F_{j,k}f\right\|_{L_{x}^{\infty}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{r}(\mathbb{R}^{2})}.$
Therefore, we reduce our proof to obtaining a decay bound at $(\frac{1}{4},\frac{1}{4})$. For any $r\in[1,\infty]$, by Hölder’s inequality, it is easy to see that
$\left\|F_{j,k}f\right\|_{L_{x}^{4}L_{u}^{r}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|F_{j,k}f\right\|_{L_{x}^{4}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}.$
Then, we just need to prove
(3.8) $\left\|F_{j,k}f\right\|_{L_{x}^{4}L_{u}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\left(-\frac{1}{4}+\epsilon\right)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$. It is easy to see that the desired decay bound at $(\frac{1}{4},\frac{1}{4})$ can be obtained by letting $\epsilon\in(0,\frac{1}{4})$ in (3.8).
To obtain (3.8), we first establish the space-time estimates for Fourier integral operators, defined on the space of Schwartz functions $\mathcal{S}(\mathbb{R}^{n})$, of the form
(3.9) $\mathcal{F}f(z):=\int_{\mathbb{R}^{n}}e^{i\phi(z,\xi)}a(z,\xi)\hat{f}(\xi)\,\mathrm{d}\xi,$
where $a(z,\xi)$ is a symbol of order $\sigma\in\mathbb{R}$. Here and hereafter, $z$ denotes a vector in $\mathbb{R}^{n}\times\mathbb{R}$ comprising the space-time variables $(x,u)$. Assume that $\textrm{supp}~a(\cdot,\xi)$ is contained in a fixed compact set and $\phi(z,\cdot)$ is homogeneous of degree 1. We say that $\mathcal{F}$ satisfies the cinematic curvature condition if the following hold:
  1. $\blacksquare$
    Non-degeneracy condition: For all $(z,\xi)\in\textrm{supp}~a$,
    $\textrm{rank}~\partial^{2}_{z\xi}\phi(z,\xi)=n.$
  2. $\blacksquare$
    Cone condition: Consider the Gauss map $G:\textrm{supp}~a\rightarrow\mathbb{S}^{n}$ defined by
    $G(z,\xi):=\frac{G_{0}(z,\xi)}{|G_{0}(z,\xi)|},\quad\text{where}\quad G_{0}(z,\xi):=\bigwedge_{j=1}^{n}\partial_{\xi_{j}}\partial_{z}\phi(z,\xi).$
    The curvature condition
    $\textrm{rank}~\partial^{2}_{\xi\xi}\langle\partial_{z}\phi(z,\xi),G(z,\xi_{0})\rangle\big|_{\xi=\xi_{0}}=n-1$
    holds for all $(z,\xi_{0})\in\textrm{supp}~a$.
We now establish the following two lemmas, which will be used repeatedly in the remaining parts of this paper.
Lemma 3.3.
Let $1\leq p\leq q\leq\infty$ and $s\in\mathbb{R}$. Let $P_{k}$ be the Littlewood-Paley operator defined in Section $1$, and $\mathcal{F}$ be the Fourier integral operator defined in (3.9) with a symbol of order $\sigma$ satisfying the cinematic curvature condition, together with the special assumption that $\textrm{rank}~\partial^{2}_{\xi\xi}\phi(z,\xi)=n-1$ for $\xi\neq 0$. Then, there exists a positive constant $C_{(u)}$ such that
$\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{q}(\mathbb{R}^{n})}\leq C_{(u)}2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{n})},$
where $s\geq\bar{s}_{n,\sigma}(p,q)$ and $\bar{s}_{n,\sigma}(p,q)$ is defined as
$\bar{s}_{n,\sigma}(p,q):=\begin{cases}\frac{1}{p}-\frac{n}{q}+\frac{n-1}{2}+\sigma,&\text{for }q>p^{\prime};\\ \frac{n}{p}-\frac{1}{q}-\frac{n-1}{2}+\sigma,&\text{for }q\leq p^{\prime}.\end{cases}$
Proof.
Indeed, this lemma can be followed from [31, Proposition 3.2] by interpolation. Here, we give another proof, which can be also applied to some other estimates for $\mathcal{F}$.
The case $q=p\in(1,\infty)$ has been proved in Seeger, Sogge and Stein [42, Theorem 2.1]. Then, by interpolation, it suffices to prove
(3.10) $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{\infty}(\mathbb{R}^{n})}\leq C_{(u)}2^{(\frac{n+1}{2}+\sigma)k}\left\|f\right\|_{L_{x}^{1}(\mathbb{R}^{n})},$
(3.11) $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{\infty}(\mathbb{R}^{n})}\leq C_{(u)}2^{(\frac{n-1}{2}+\sigma)k}\left\|f\right\|_{L_{x}^{\infty}(\mathbb{R}^{n})},$
and
(3.12) $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{1}(\mathbb{R}^{n})}\leq C_{(u)}2^{(\frac{n-1}{2}+\sigma)k}\left\|f\right\|_{L_{x}^{1}(\mathbb{R}^{n})}.$
To establish (3.10), by changing variables, we express the kernel of $\mathcal{F}P_{k}$ as
$K_{k}(z,y):=2^{nk}\int_{\mathbb{R}^{n}}e^{i2^{k}(\phi(z,\xi)-y\cdot\xi)}a(z,2^{k}\xi)\bar{\psi}(|\xi|)\,\textrm{d}\xi.$
Without loss of generality, we may suppose that $\textrm{rank}~\partial_{\xi^{\prime}\xi^{\prime}}^{2}\phi(z,\xi)=n-1$ where $\xi:=(\xi_{1},\xi^{\prime})$ with $\xi^{\prime}:=(\xi_{2},\cdots,\xi_{n})$ is perpendicular to $\xi_{1}$. By partition of unity, we can further suppose that $\bar{\psi}=0$ if $|\frac{\xi}{|\xi|}-e_{1}|\geq\frac{1}{10^{2}}$ and, for all $z\in\textrm{supp}~a(\cdot,\xi)$ and $y\in\mathbb{R}^{n}$, there is at most one $v\in\mathbb{S}^{n-1}$ satisfying $|v-e_{1}|\leq\frac{1}{9^{2}}$ such that
(3.13) $\nabla_{\xi}\phi(z,v)=y,$
where $e_{1}:=(1,0,\cdots,0)\in\mathbb{S}^{n-1}$. If such $v$ does not exist, then the phase function in $K_{k}$ has no critical point. Moreover, by homogeneity, it follows that
$|K_{k}(z,y)|\leq C_{(u,N)}2^{-Nk}$
if $N$ is large enough, which further leads to (3.10).
Now, we suppose that there exists such $v$. Without loss of generality we may assume that $v=e_{1}$, otherwise we can choose new axes so that $v=e_{1}$. Since $\phi(z,\xi)$ is homogeneous of degree $1$ in $\xi$,
$\nabla_{\xi}\phi(z,\xi)\cdot\xi=\phi(z,\xi);$
hence $\nabla_{\xi}(\partial_{\xi_{i}}\phi)(z,\xi)\cdot\xi=0$ for any $i=1,2,\cdots,n$, furthermore, $\partial^{2}_{\xi_{i}\xi_{1}}\phi(z,e_{1})=0$ holds for all $i=1,2,\cdots,n$. This, combined with he special assumption that $\textrm{rank}~\partial^{2}_{\xi^{\prime}\xi^{\prime}}\phi(z,\xi)=n-1$ when $\xi\neq 0$, leads to
(3.14) $\det\partial_{\xi^{\prime}\xi^{\prime}}^{2}\phi(z,e_{1})\neq 0.$
It is easy to see that, for any $i,j=1,2,\cdots,n$, $\partial^{2}_{\xi_{i}\xi_{j}}\phi(z,\xi)$ is homogeneous of degree $1$ in $\xi$ and $\partial_{\xi_{i}}\phi(z,\xi)$ is homogeneous of degree 0 in $\xi$. Let $\phi_{\xi_{1}}(z,\xi^{\prime}):=\phi(z,\xi)$. Then, for all $|(\xi_{1},0)|\in\textrm{supp}~\bar{\psi}$, by (3.13) and (3.14), we can see that
$\nabla_{\xi^{\prime}}\phi_{\xi_{1}}(z,0)=\nabla_{\xi^{\prime}}\phi(z,(\xi_{1},0))=\nabla_{\xi^{\prime}}\phi(z,e_{1})=y^{\prime},$
and
$\det\partial_{\xi^{\prime}\xi^{\prime}}^{2}\phi_{\xi_{1}}(z,0)=\det\partial_{\xi^{\prime}\xi^{\prime}}^{2}\phi(z,(\xi_{1},0))=|\xi_{1}|^{-(n-1)}\det\partial_{\xi^{\prime}\xi^{\prime}}^{2}\phi(z,e_{1})\neq 0,$
where $y:=(y_{1},y^{\prime})$. Therefore, from Stein [47, Chapter VIII, Proposition 6], there exists a constant $C$ such that the following identity is valid:
(3.15) $K_{k}(z,y)$ $=2^{nk}\int_{\mathbb{R}}\int_{\mathbb{R}^{n-1}}e^{i2^{k}(\phi(z,\xi)-y\cdot\xi)}a(z,2^{k}\xi)\bar{\psi}(|\xi|)\,\textrm{d}\xi^{\prime}\,\textrm{d}\xi_{1}$
$=C2^{\frac{n+1}{2}k}\int_{\mathbb{R}}e^{i2^{k}\left(\phi(z,(\xi_{1},0))-y_{1}\xi_{1}\right)}a\left(z,(2^{k}\xi_{1},0)\right)\bar{\psi}(|\xi_{1}|)\,\textrm{d}\xi_{1}+O\left(2^{\left(\frac{n-1}{2}+\sigma\right)k}\right),$
which further leads to $|K_{k}(z,y)|\lesssim 2^{\left(\frac{n+1}{2}+\sigma\right)k}$ and (3.10) follows easily.
We now turn to the proof of (3.11), which is based on some results obtained in Stein [47, Chapter IX, Section 4]. Let $\mathcal{F}$ be the Fourier integral operator with a symbol of order $\frac{n-1}{2}$, it is enough to prove
$\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{\infty}(\mathbb{R}^{n})}\leq C_{(u)}\left\|f\right\|_{L_{x}^{\infty}(\mathbb{R}^{n})},$
which can be reduced to establish
(3.16) $\left|\int_{\mathbb{R}^{n}}K_{k}(z,y)\,\textrm{d}y\right|\leq C_{(u)},$
where the kernel of $\mathcal{F}P_{k}$ is given by
$K_{k}(z,y)=\int_{\mathbb{R}^{n}}e^{i(\phi(z,\xi)-y\cdot\xi)}a(z,\xi)\bar{\psi}_{k}(|\xi|)\,\textrm{d}\xi.$
Next, a further dyadic division is needed. Roughly speaking, we will split each dyadic shell $\{\xi\in\mathbb{R}^{n}:2^{k-3}\leq|\xi|\leq 2^{k+3}\}$ into thin truncated cones of aperture roughly $2^{-\frac{k}{2}}$, and each such truncated cone is essentially an elongated rectangle whose major axis has length roughly $2^{k}$, while all the other sides have length roughly $2^{\frac{k}{2}}$. Then, we need roughly $2^{\frac{n-1}{2}k}$ such truncated cones to cover the shell $\{\xi\in\mathbb{R}^{n}:2^{k-3}\leq|\xi|\leq 2^{k+3}\}$ (see Stein [47, Chapter IX, Figure 1]). More precisely, for each $k\in\mathbb{N}$, we fix a collection $\{\xi_{k}^{\nu}\}_{\nu}\subset\mathbb{S}^{n-1}$ satisfy:
  1. (1)
    $|\xi_{k}^{\nu}-\xi_{k}^{\nu^{\prime}}|\geq 2^{-\frac{k}{2}}$, if $\nu\neq\nu^{\prime}$;
  2. (2)
    If $\xi\in\mathbb{S}^{n-1}$, then there exists a $\xi_{k}^{\nu}$ so that $|\xi-\xi_{k}^{\nu}|<2^{-\frac{k}{2}}$.
Let
$\Gamma_{k}^{\nu}:=\left\{\xi\in\mathbb{R}^{n}:\left|\frac{\xi}{|\xi|}-\xi_{k}^{\nu}\right|\leq 2^{1-\frac{k}{2}}\right\},$
which is a cone with central direction $\xi_{k}^{\nu}$. We can then construct an associated partition of unity
$\sum_{\nu}\chi_{k}^{\nu}(\xi)=1$
for all $\xi\neq 0$ and $k\in\mathbb{N}$, where $\chi_{k}^{\nu}$ is homogeneous of degree 0 in $\xi$, supported in $\Gamma_{k}^{\nu}$ and satisfy $|\partial_{\xi}^{\alpha}\chi_{k}^{\nu}(\xi)|\leq C_{(\alpha)}2^{\frac{|\alpha|}{2}k}|\xi|^{-|\alpha|}$ for all $\alpha\in\mathbb{N}_{0}^{n}$.
By the definition of $\chi_{k}^{\nu}$, let
$K^{\nu}_{k}(z,y):=\int_{\mathbb{R}^{n}}e^{i(\phi(z,\xi)-y\cdot\xi)}a(z,\xi)\bar{\psi}_{k}(|\xi|)\chi_{k}^{\nu}(\xi)\,\textrm{d}\xi.$
Consequently,
(3.17) $K_{k}=\sum_{\nu}K^{\nu}_{k}.$
Furthermore, we choose axes in the $\xi$-space so that $\xi_{1}$ is in the direction of $\xi_{k}^{\nu}$ and write
$K^{\nu}_{k}(z,y)=\int_{\mathbb{R}^{n}}e^{i(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y)\cdot\xi}b_{k}^{\nu}(z,\xi)\,\textrm{d}\xi,$
where
$b_{k}^{\nu}(z,\xi):=e^{i(\phi(z,\xi)-\nabla_{\xi}\phi(z,\xi_{k}^{\nu})\cdot\xi)}a(z,\xi)\bar{\psi}_{k}(|\xi|)\chi_{k}^{\nu}(\xi).$
We define the following operator
$L:=I-2^{2k}\partial^{2}_{\xi_{1}\xi_{1}}-2^{k}\nabla_{\xi^{\prime}}.$
As in Stein [47, Chapter IX, Section 4], for any $N\in\mathbb{N}_{0}$, we have
$L^{N}\left(e^{i(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y)\cdot\xi}\right)=\left[1+2^{2k}\left|\left(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y\right)_{1}\right|^{2}+2^{k}\left|\left(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y\right)^{\prime}\right|^{2}\right]^{N}e^{i(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y)\cdot\xi},$
where $(\cdot)_{1}$ denotes the component in the direction $\xi_{k}^{\nu}$ and $(\cdot)^{\prime}$ denotes the orthogonal component. Moreover, by the fact that $\mathcal{F}$ has a symbol of order $\frac{n-1}{2}$, it follows that
$L^{N}\left(b_{k}^{\nu}(z,\xi)\right)\leq C_{(u,N)}2^{-\frac{n-1}{2}k}.$
Notice that the support of $b_{k}^{\nu}(z,\cdot)$ has volume roughly $2^{(1+\frac{n-1}{2})k}$, pass the operator $L$ onto $K^{\nu}_{k}$, we deduce that
(3.18) $\left|K^{\nu}_{k}(z,y)\right|\leq C_{(u,N)}2^{k}\left[1+2^{k}\left|\left(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y\right)_{1}\right|+2^{\frac{k}{2}}\left|\left(\nabla_{\xi}\phi(z,\xi_{k}^{\nu})-y\right)^{\prime}\right|\right]^{-2N}.$
If we take $N$ large enough, it is easy to see that
$\left|\int_{\mathbb{R}^{n}}K^{\nu}_{k}(z,y)\,\textrm{d}y\right|\leq C_{(u)}2^{-\frac{n-1}{2}k}.$
This, combined with (3.17) and the fact that there are essentially $2^{\frac{n-1}{2}k}$ terms involved, leads to (3.16), as desired.
The proof of (3.12) is very similar to that of (3.11). Indeed, the change of variables $x\rightarrow\nabla_{\xi}\phi(z,\xi_{k}^{\nu})$ in (3.18), whose Jacobian is bounded from below, leads to
$\left|\int_{\mathbb{R}^{n}}K^{\nu}_{k}(z,y)\,\textrm{d}x\right|\leq C_{(u)}2^{-\frac{n-1}{2}k}.$
Then,
$\left|\int_{\mathbb{R}^{n}}K_{k}(z,y)\,\textrm{d}x\right|\leq C_{(u)},$
which implies (3.12). The proof is thus complete. ∎
Lemma 3.4.
Let $n=2$, $1\leq p\leq q\leq\infty$, and $s\in\mathbb{R}$. Let $P_{k}$ be the Littlewood-Paley operator defined in Section $1$, and $\mathcal{F}$ be a Fourier integral operator as in (3.9) with a symbol of order $\sigma$ satisfying the cinematic curvature condition, and assume $\textrm{rank}~\partial^{2}_{\xi\xi}\phi(z,\xi)=1$ for $\xi\neq 0$. Then, for $P_{k}$ defined as in Section 1 with $k\in\mathbb{N}$, there exists a constant $C>0$ such that
$\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times[1,2])}\leq C2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $s\geq s_{\sigma,\epsilon}(p,q,q)$ and, for any $\epsilon>0$,
$s_{\sigma,\epsilon}(p,q,q):=\begin{cases}\frac{1}{p}-\frac{3}{q}+\frac{1}{2}+\sigma+\epsilon,&\text{for }q\geq 3p^{\prime};\\ \frac{3}{2p}-\frac{3}{2q}+\sigma+\epsilon,&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-\frac{1}{q}-\frac{1}{2}+\sigma,&\text{for }q\leq p^{\prime}.\end{cases}$
Proof.
For any $\epsilon>0$, by interpolation, the key estimates we shall make are:
  1. (a)
    $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{\infty}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{\left(\frac{3}{2}+\sigma+\epsilon\right)k}\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}$;
  2. (b)
    $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{\infty}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{\left(\frac{1}{2}+\sigma+\epsilon\right)k}\|f\|_{L_{x}^{\infty}(\mathbb{R}^{2})}$;
  3. (c)
    $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{1}L^{1}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{\left(\frac{1}{2}+\sigma\right)k}\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}$;
  4. (d)
    $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{2}L^{2}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{\sigma k}\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}$;
  5. (e)
    $\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(\sigma+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$.
The proofs of estimates $(\textrm{a})-(\textrm{d})$ follow directly from Lemma 3.3, since $C_{(u)}\lesssim 1$ for any $u\in[1,2]$. On the other hand, we want to show that Lee [26, Corollary 1.5] also implies the estimates $(\textrm{a})-(\textrm{b})$, where the following additional condition is needed:
all nonzero eigenvalues of $\partial^{2}_{\xi\xi}\langle\partial_{z}\phi(z,\xi),G(z,\xi_{0})\rangle\big|_{\xi=\xi_{0}}$ have the same sign.
However, this additional condition is clearly satisfied, since the cone condition
$\textrm{rank}~\partial^{2}_{\xi\xi}\langle\partial_{z}\phi(z,\xi),G(z,\xi_{0})\rangle\big|_{\xi=\xi_{0}}=1$
holds for all $(z,\xi_{0})\in\textrm{supp}~a$ when $n=2$, which further implies that there is only one nonzero eigenvalue of $\partial^{2}_{\xi\xi}\langle\partial_{z}\phi(z,\xi),G(z,\xi_{0})\rangle\big|_{\xi=\xi_{0}}$, and so has the same sign.
The estimate $(\textrm{c})$ can also be proved in the similar way as (3.10). Indeed, by the fact that $\phi(z,\xi)$ is homogeneous of degree $1$ in $\xi$, we can write (3.15) as
$K_{k}(z,y)=C2^{\frac{3}{2}k}\int_{\mathbb{R}}e^{i2^{k}\left(\phi(z,e_{1})-y_{1}\right)\xi_{1}}a\left(z,(2^{k}\xi_{1},0)\right)\bar{\psi}(|\xi_{1}|)\,\textrm{d}\xi_{1}+O\left(2^{\left(\frac{1}{2}+\sigma\right)k}\right),$
when $n=2$. Then, interpolation by parts shows that
$|K_{k}(z,y)|\lesssim 2^{\left(\frac{3}{2}+\sigma\right)k}(1+2^{k}|\phi(z,e_{1})-y_{1}|)^{-N}+2^{\left(\frac{1}{2}+\sigma\right)k},$
where $N\in\mathbb{N}$ is large enough. Furthermore, from Mockenhaupt, Seeger and Sogge [36, P. 70], by the above non-degeneracy condition, we can choose a locally coordinates $z=(x,u)\in\mathbb{R}^{2}\times\mathbb{R}$ such that
$\textrm{rank}~\partial^{2}_{x\xi}\phi(z,\xi)=2,$
and, if $\xi\neq 0$,
$\partial_{u}\phi(z,\xi)\neq 0.$
It follows that $|\partial_{u}\phi(z,e_{1})|\gtrsim 1$ for all $z\in\textrm{supp}~a(\cdot,\xi)$. Therefore, via a change of variables $u\rightarrow\phi(z,e_{1})$, for any $z\in\textrm{supp}~a(\cdot,\xi)$, we assert that
$\int_{\mathbb{R}}|K_{k}(z,y)|\,\textrm{d}u$ $\lesssim 2^{\left(\frac{3}{2}+\sigma\right)k}\int_{\mathbb{R}}(1+2^{k}|\phi(z,e_{1})-y_{1}|)^{-N}\,\textrm{d}u+2^{\left(\frac{1}{2}+\sigma\right)k}$
$\lesssim 2^{\left(\frac{3}{2}+\sigma\right)k}\int_{\mathbb{R}}(1+2^{k}|u-y_{1}|)^{-N}\,\textrm{d}u+2^{\left(\frac{1}{2}+\sigma\right)k}\lesssim 2^{\left(\frac{1}{2}+\sigma\right)k}.$
This, combined with the fact that $\textrm{supp}~a(\cdot,\xi)$ is contained in a fixed compact set, leads to the estimate $(\textrm{c})$.
The estimate $(\textrm{d})$ is a basic result about the Fourier integral operators $\mathcal{F}$, which can be proved by considering the pseudo-differential operator $\mathcal{F}^{\ast}\mathcal{F}$, we refer the reader to $\textrm{H}\ddot{\textrm{o}}\textrm{rmander}$ [20], Stein [47, Chapter IX, Section 3] or Sogge [43, Section 6.2] for more detailed discussions. The key estimate $(\textrm{e})$ follows from Gao, Liu, Miao and Xi [12, Theorem 1.4], which is a generalization of the local smoothing conjecture in $2+1$ dimensions obtained in Guth, Wang and Zhang [16, Theorem 1.2]. This finishes the proof. ∎
Remark 3.5.
From [31], we can see that the result in Lemma 3.4 is nearly sharp. Moreover, if we restrict our attention to the Littlewood-Paley operator $P_{k}$, Lemma 3.4 extends the $L_{x}^{2}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate obtained in Mockenhaupt, Seeger and Sogge [36, Theorem 3.1] or the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate with $q\geq 3p^{\prime}$ and $q\geq\frac{5}{3}p$ proved in Lee [26, Corollary 1.5] to an almost sharp $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate. Specifically, we have
$\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
with
$\begin{cases}\sigma<-\frac{1}{p}+\frac{3}{q}-\frac{1}{2},&\text{for }q\geq 3p^{\prime};\\ \sigma<-\frac{3}{2p}+\frac{3}{2q},&\text{for }p^{\prime}<q<3p^{\prime};\\ \sigma\leq-\frac{2}{p}+\frac{1}{q}+\frac{1}{2},&\text{for }q\leq p^{\prime}.\end{cases}$
Remark 3.6.
In [31], for $F_{j,k}$ defined in (2.9), we established
(3.19) $\left\|F_{j,k}f\right\|_{L_{x}^{6}L^{6}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\epsilon k}\|f\|_{L_{x}^{2}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$. This estimate can be derived by interpolating between (3.2) and the $F_{j,k}:L_{x}^{4}(\mathbb{R}^{2})\rightarrow L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate, the latter follows from Gao, Liu, Miao and Xi [12, Theorem 1.4]. Indeed, (3.19) can be improved to
(3.20) $\left\|F_{j,k}f\right\|_{L_{x}^{6}L^{6}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{2}(\mathbb{R}^{2})},$
which follows from Seeger, Sogge and Stein [42, Theorem 3.1] by the fact that $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ satisfying the cinematic curvature condition. We remark that the non-degeneracy assumptions (2.2)–(2.3) and the cone condition (2.6) in Mockenhaupt, Seeger and Sogge [36, Theorem 3.1] are equivalent to the cinematic curvature condition.
On the other hand, we may see from [31] that the estimate (3.20) is an endpoint estimate, which implies the sharpness of the $L^{2}_{\textrm{comp}}(Y)\rightarrow L^{q}_{\textrm{loc}}(Z)$ local smoothing theorem [36, Theorem 3.1], and further implies the sharpness of the square function estimate [36, Theorem 3.2] by using Sobolev’s embedding theorem in the time variable $u$, at least in two-dimensional situation. Here, we also note that the results in [36, Theorems 3.1 and 3.2] actually are sharp for any $n$-dimensional setting by using the counterexample in Sogge [43, Section 6.2, P. 185].
Now, we turn to the proof of (3.8). Notice that $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ and satisfies (2.10), then the estimate $(\textrm{e})$ in Lemma 3.4 implies
(3.21) $\left\|F_{j,k}f\right\|_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(-\frac{1}{2}+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$. Similarly, we can obtain that
(3.22) $\left\|\partial_{u}(F_{j,k}f)\right\|_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(\frac{1}{2}+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}.$
Therefore, as an immediate consequence of (3.21) and (3.22), by Lemma 3.2, we deduce that
$\left\|F_{j,k}f\right\|_{L_{x}^{4}L^{\infty}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim\left\|F_{j,k}f\right\|^{\frac{3}{4}}_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\left\|\partial_{u}(F_{j,k}f)\right\|^{\frac{1}{4}}_{L_{x}^{4}L^{4}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim 2^{(-\frac{1}{4}+\epsilon)k}\|f\|_{L_{x}^{4}(\mathbb{R}^{2})}$
for any $\epsilon\in(0,\infty)$, which is (3.8) as desired.
Putting things together, we complete the proof of the sufficiency of the region $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$.
3.2 The necessity of the region of $(\frac{1}{p},\frac{1}{q})$ for all $r\in[1,\infty]$
In this subsection, we establish the necessary conditions for Theorem 1.1. The condition $0\leq\frac{1}{q}$ is trivial, and the condition $\frac{1}{q}\leq\frac{1}{p}$ is a standard necessity following from a theorem of Hörmander [19].
We now demonstrate that, if
(3.23) $\left\|Tf\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
holds, then $(\frac{1}{p},\frac{1}{q})$ satisfies
$1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)\geq 0.$
The proof is similar to [31]. Specifically, it suffices to show
(3.24) $t_{i}\left(t_{i}\left|\gamma(t_{i})\right|\right)^{\frac{1}{q}-\frac{1}{p}}\lesssim 1,$
where the implicit constant is independent of $i$, and $\{t_{i}\}_{i=1}^{\infty}$ strictly decreasing and converging to 0. Indeed, let $S_{t_{i}}:=[-t_{i},t_{i}]\times[-|\gamma(t_{i})|,|\gamma(t_{i})|]$. As shown in [31] that
$\delta t_{i}\chi_{\frac{S_{t_{i}}}{2}}(x)\leq T\chi_{S_{t_{i}}}(x,u)$
holds for some $\delta$ sufficiently small. Combining this with Hölder’s inequality and (3.23), we obtain
$\int_{1}^{2}\left|\frac{S_{t_{i}}}{2}\right|\,\mathrm{d}u\leq\frac{1}{\delta t_{i}}\int_{\frac{S_{t_{i}}}{2}}\left\|T\chi_{S_{t_{i}}}(x,u)\right\|_{L^{1}_{u}([1,2])}\,\mathrm{d}x\lesssim\frac{|S_{t_{i}}|^{\frac{1}{q^{\prime}}}}{t_{i}}\left\|T\chi_{S_{t_{i}}}\right\|_{L_{x}^{q}L^{1}_{u}(\mathbb{R}^{2}\times[1,2])}\lesssim\frac{|S_{t_{i}}|^{\frac{1}{q^{\prime}}}}{t_{i}}\left\|\chi_{S_{t_{i}}}\right\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Since $|S_{t_{i}}|\approx t_{i}|\gamma(t_{i})|$ and $\|\chi_{S_{t_{i}}}\|_{L_{x}^{p}(\mathbb{R}^{2})}\approx(t_{i}|\gamma(t_{i})|)^{\frac{1}{p}}$, the desired inequality (3.24) follows.
To obtain the other necessary conditions for Theorem 1.1, we also require the following proposition, which is based on Lemmas 3.2, 3.3 and 3.4.
Proposition 3.7.
Let $n=2$, $1\leq p\leq q\leq\infty$, and $\mathcal{F}$ be as in Lemma 3.4. Then, there exists a constant $C>0$ such that
$\left\|\mathcal{F}P_{k}f\right\|_{L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\leq C2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $s\geq s_{\sigma,\epsilon}(p,q,\infty)$ and, for any $\epsilon>0$,
$s_{\sigma,\epsilon}(p,q,\infty):=\begin{cases}\frac{1}{p}-\frac{2}{q}+\frac{1}{2}+\sigma+\epsilon,&\text{for }q\geq 3p^{\prime};\\ \frac{3}{2p}-\frac{1}{2q}+\sigma+\epsilon,&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-\frac{1}{2}+\sigma,&\text{for }q\leq p^{\prime}.\end{cases}$
Proof.
Notice that $\phi(z,\cdot)$ is a homogeneous function of degree $1$, which leads to that $\partial_{u}\phi(z,\cdot)$ is also a homogeneous function of degree $1$ and $|\partial_{u}\phi(z,\cdot)|\lesssim|\xi|$. Then, the proof of the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L_{x}^{q}L^{q}_{u}(\mathbb{R}^{2}\times\mathbb{R})$ estimate for $\partial_{u}(\mathcal{F}P_{k}f)$ is very similar as that for $\mathcal{F}P_{k}f$, but need to multiply a extra factor $2^{k}$. Therefore, by Lemma 3.2 with $r=q$ and Lemma 3.4, we can say
$s_{\sigma,\epsilon}(p,q,\infty)=s_{\sigma,\epsilon}(p,q,q)+\frac{1}{q}$
and complete the proof of this proposition. ∎
Let $Q_{k}f:=(\varrho_{k}(\xi)\hat{f}(\xi))^{\vee}$ with $\varrho_{k}(\xi):=\varrho(2^{-k}\xi)$ and $k\in\mathbb{N}$, where $\varrho\in\mathcal{S}(\mathbb{R}^{2})$ and
$\textrm{supp}~\varrho\subset\left\{\xi\in\mathbb{R}^{2}:\frac{1}{2}\leq|\xi|\leq 1\text{ and }\frac{|\xi_{1}|}{|\xi_{2}|}\in\textrm{supp}~\rho\right\}.$
Assume that (3.23) holds for some $1\leq p\leq q\leq\infty$ and $1\leq r\leq\infty$, and apply it to $Q_{k}f$. From the preliminaries in Section 1, we conclude that
$\left\|F_{j,k}P_{k}Q_{k}f\right\|_{L_{x}^{q}L^{r}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Here, we used the fact that $F_{j,k}P_{k}Q_{k}f=F_{j,k}Q_{k}f$. Furthermore, by Lemma 3.2, we have
(3.25) $\left\|F_{j,k}P_{k}Q_{k}f\right\|_{L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\frac{1}{r}k}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
On the other hand, since $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ satisfying the special assumption (2.10), Proposition 3.7 implies
(3.26) $\left\|F_{j,k}P_{k}Q_{k}f\right\|_{L_{x}^{q}L^{\infty}_{u}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
with $s\geq s_{-\frac{1}{2},\epsilon}(p,q,\infty)$, where
$s_{-\frac{1}{2},\epsilon}(p,q,\infty)=\begin{cases}\frac{1}{p}-\frac{2}{q}+\epsilon,&\text{for }q\geq 3p^{\prime};\\ \frac{3}{2p}-\frac{1}{2q}-\frac{1}{2}+\epsilon,&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-1,&\text{for }q\leq p^{\prime}.\end{cases}$
Here $\epsilon\in(0,\infty)$ is an arbitrary constant. Moreover, from [35], the estimates (3.1), (3.5), and (3.7) represent the endpoint estimates for $F_{j,k}$ when taking the $L^{\infty}$-norm in $u$. i.e., the points $(0,0)$, $(\frac{1}{2},\frac{1}{2})$, and $(\frac{2}{5},\frac{1}{5})$ are the endpoints. Consequently, $s_{-\frac{1}{2},0}(p,q,\infty)$ is the infimum such that (3.26) holds. Combining this with (3.25), we obtain $\frac{1}{r}\geq\max\{\frac{1}{p}-\frac{2}{q},\frac{3}{2p}-\frac{1}{2q}-\frac{1}{2},\frac{2}{p}-1\}$. This further implies
$\max\left\{\frac{1}{2p}-\frac{1}{2r},\frac{3}{p}-\frac{r+2}{r}\right\}\leq\frac{1}{q}~~\textrm{and}~~\frac{1}{p}\leq\frac{r+1}{2r},$
as desired.
Therefore, we conclude the proof of the necessity of $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.1 for all $r\in[1,\infty]$.
4 Proof of Theorem 1.4
In this section, we investigate the time-space estimate, i.e., the $L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])$ estimate, for the operator $T$ defined in (2.1). We address the sufficient and necessary parts separately, although the result is not sharp due to certain technical limitations.
4.1 Sufficient Part
The proof of the time-space estimate for $T$ differs significantly from the space-time estimate established in Theorem 1.1. From the preliminaries in Section 2, it suffices to prove that
$\sum_{j\in\mathbb{Z}_{0}^{-}}2^{j}|2^{j}\gamma(2^{j})|^{\frac{1}{q}-\frac{1}{p}}\left\|\widetilde{T}_{j}\right\|_{L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim 1.$
By Lemma 2.3, it remains to prove the time-space estimate for $\widetilde{T}_{j}$ uniformly in $j\in\mathbb{Z}_{0}^{-}$. As in Section 2, we reduce this estimate to $\widetilde{T}^{1}_{j,k}$ defined in (2.3). Furthermore, from [35], we know that the time-space estimate for $\widetilde{T}^{1,a}_{j,k}$ is bounded by $2^{-Nk}$ for $N$ sufficiently large. We can then reduce to estimate the operator $\widetilde{T}^{1,b}_{j,k}$, rather than $F_{j,k}$, in the case $1\leq r<p\leq q\leq\infty$.
Indeed, for $F_{j,k}$ and $E_{j,k}$ defined in (2.6) and (2.7), respectively, we also decompose $\mathbb{R}^{2}$ into $\bigcup_{i\in\mathbb{Z}^{2}}B(x_{i},\varpi)$ such that $|x_{i}-x_{i^{\prime}}|\approx|i-i^{\prime}|\varpi$ for any $i\neq i^{\prime}$. Unlike (2.8), we bound $\|\widetilde{T}^{1,b}_{j,k}f\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}$ by
(4.1) $\left\|\left[\sum_{i\in\mathbb{Z}^{2}}\left\|\left(E_{j,k}-F_{j,k}\right)\left(f(\cdot-x_{i})\right)\right\|^{q}_{L_{x}^{q}(\mathbb{R}^{2})}\right]^{\frac{1}{q}}\right\|_{L_{u}^{r}(\mathbb{R})}+\left\|\left[\sum_{i\in\mathbb{Z}^{2}}\left\|F_{j,k}\left(f(\cdot-x_{i})\right)\right\|^{q}_{L_{x}^{q}(\mathbb{R}^{2})}\right]^{\frac{1}{q}}\right\|_{L_{u}^{r}(\mathbb{R})}.$
The estimate for the first term follows from an argument similar to that for $\widetilde{T}^{1,a}_{j,k}$. The key difference lies in the second term. To be more precise, when $1\leq p\leq q\leq r\leq\infty$, Minkowski’s inequality implies that the second term can be bounded by
$\left[\sum_{i\in\mathbb{Z}^{2}}\left\|F_{j,k}\left(f(\cdot-x_{i})\right)\right\|^{q}_{L_{u}^{r}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\right]^{\frac{1}{q}},$
which can be reduced to the corresponding time-space estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\rightarrow L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ with a decay in $k$. When $1\leq p\leq r\leq q\leq\infty$, the second term can be bounded as
$\left\|\left[\sum_{i\in\mathbb{Z}^{2}}\left\|F_{j,k}\left(f(\cdot-x_{i})\right)\right\|^{r}_{L_{x}^{q}(\mathbb{R}^{2})}\right]^{\frac{1}{r}}\right\|_{L_{u}^{r}(\mathbb{R})}\leq\left[\sum_{i\in\mathbb{Z}^{2}}\left\|F_{j,k}\left(f(\cdot-x_{i})\right)\right\|^{p}_{L_{u}^{r}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\right]^{\frac{1}{p}},$
which also reduces to the corresponding decay time-space estimate for $F_{j,k}$. Thus, the only case where we do not reduce the estimate to a corresponding decay time-space estimate for $F_{j,k}$ is when $1\leq r<p\leq q\leq\infty$.
We remark that the most challenging aspect of proving the time-space estimates for $T$ may not lie in the case $1\leq r<p\leq q\leq\infty$, since the $L_{x}^{\infty}(\mathbb{R}^{2})\rightarrow L^{\infty}_{u}L_{x}^{\infty}(\mathbb{R}^{2}\times[1,2])$ estimate for $T$ holds trivially, and we can handle the region $1\leq r<p\leq q\leq\infty$ via interpolation. The most significant difference is that we can only obtain a common estimate
$\|F_{j,k}f\|_{L^{r}_{u}L_{x}^{\infty}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{k}\|f\|_{L_{x}^{1}(\mathbb{R}^{2})}$
for all $r\in[1,\infty]$, unlike the estimate in (3.3).
Based on the above discussion, we divide the proof of Theorem 1.4 into three parts.
4.1.1 The Case $1\leq p\leq q\leq r\leq\infty$
Observe that $F_{j,k}P_{k}f=F_{j,k}f$, and $F_{j,k}$ is a Fourier integral operator of order $\frac{1}{2}$ satisfying (2.10). By Lemma 3.4, we have
(4.2) $\left\|F_{j,k}f\right\|_{L^{q}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $s\geq s_{-\frac{1}{2},\epsilon}(p,q,q)$ and
$s_{-\frac{1}{2},\epsilon}(p,q,q)=\begin{cases}\frac{1}{p}-\frac{3}{q}+\epsilon,&\text{for }q\geq 3p^{\prime};\\ \frac{3}{2p}-\frac{3}{2q}-\frac{1}{2}+\epsilon,&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-\frac{1}{q}-1,&\text{for }q\leq p^{\prime}.\end{cases}$
Additionally, Lemma 3.3 implies [^2] We remark that: by the following Lemma 4.2, then the estimate (4.3) is established with $s\geq s_{-\frac{1}{2},\epsilon}(p,q,q)+\frac{1}{q}$. However, we have $\bar{s}_{2,-\frac{1}{2}}(p,q)<s_{-\frac{1}{2},\epsilon}(p,q,q)+\frac{1}{q}$.
(4.3) $\left\|F_{j,k}f\right\|_{L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $s\geq\bar{s}_{2,-\frac{1}{2}}(p,q)$ with
$\bar{s}_{2,-\frac{1}{2}}(p,q):=\begin{cases}\frac{1}{p}-\frac{2}{q},&\text{for }q>p^{\prime};\\ \frac{2}{p}-\frac{1}{q}-1,&\text{for }q\leq p^{\prime}.\end{cases}$
Interpolating between (4.2) and (4.3), we obtain
(4.4) $\left\|F_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
for all $1\leq p\leq q\leq r\leq\infty$, where $s\geq\tilde{s}_{A,\epsilon}(p,q,r)$ and, for any $\epsilon>0$,
$\tilde{s}_{A,\epsilon}(p,q,r):=\begin{cases}\frac{1}{p}-\frac{2}{q}-\frac{1}{r}+\epsilon,&\text{for }q\geq 3p^{\prime};\\ \frac{1}{p}-\frac{2}{q}+\left(\frac{1}{2p}+\frac{1}{2q}-\frac{1}{2}\right)\frac{q}{r}+\epsilon,&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{2}{p}-\frac{1}{q}-1,&\text{for }q\leq p^{\prime}.\end{cases}$
Therefore, when $(\frac{1}{p},\frac{1}{q})$ satisfies $\tilde{s}_{A,0}(p,q,r)<0$, we obtain a desired decay time-space estimate for $F_{j,k}$, which in turn implies the time-space estimate for $T$.
4.1.2 The Case $1\leq p\leq r\leq q\leq\infty$
One can apply the Sobolev embedding theorem with respect to the space variable $x$ to see that
$\left\|F_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{2\left(\frac{1}{r}-\frac{1}{q}\right)k}\left\|F_{j,k}f\right\|_{L^{r}_{u}L_{x}^{r}(\mathbb{R}^{2}\times\mathbb{R})}$
for $r\leq q\leq\infty$. Combining this with (4.4), we derive
(4.5) $\left\|F_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
for all $1\leq p\leq r\leq q\leq\infty$, where $s\geq\tilde{s}_{A,\epsilon}(p,r,r)+2\left(\frac{1}{r}-\frac{1}{q}\right)$.
On the other hand, by Hölder’s inequality and (4.4), a straightforward computation shows that
$\left\|F_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|F_{j,k}f\right\|_{L^{q}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
holds for all $1\leq p\leq r\leq q\leq\infty$ with $s\geq\tilde{s}_{A,\epsilon}(p,q,q)$. Combining this with (4.5), we obtain
(4.6) $\left\|F_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
for all $1\leq p\leq r\leq q\leq\infty$, where $s\geq\tilde{s}_{B,\epsilon}(p,q,r)$ with
$\tilde{s}_{B,\epsilon}(p,q,r):=\min\left\{\tilde{s}_{A,\epsilon}(p,r,r)+2\left(\frac{1}{r}-\frac{1}{q}\right),\tilde{s}_{A,\epsilon}(p,q,q)\right\}.$
Then, when $(\frac{1}{p},\frac{1}{q})$ satisfies $\tilde{s}_{B,0}(p,q,r)<0$, we obtain a desired decay time-space estimate for $F_{j,k}$, which implies the time-space estimate for $T$.
4.1.3 The Case $1\leq r<p\leq q\leq\infty$
As in [35], by (4.1), the corresponding time-space estimate for $E_{j,k}-F_{j,k}$ has a bound of $2^{-Nk}$ for $N$ sufficiently large, for all $1\leq p\leq q\leq\infty$ and $1\leq r\leq\infty$. Thus, the estimate in (4.5) also applies to $\widetilde{T}^{1,b}_{j,k}$. Combining this with the time-space estimate for $\widetilde{T}^{1,a}_{j,k}$, we obtain
(4.7) $\left\|\widetilde{T}^{1}_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
for all $1\leq p\leq r\leq q\leq\infty$ with $s\geq\tilde{s}_{B,\epsilon}(p,q,r)$.
Now, consider the case $1\leq r<p\leq q\leq\infty$. By Hölder’s inequality and (4.7), we have
$\left\|\widetilde{T}^{1}_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|\widetilde{T}^{1}_{j,k}f\right\|_{L^{p}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\tilde{s}_{B,\epsilon}(p,q,p)k}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
and
$\left\|\widetilde{T}^{1}_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|\widetilde{T}^{1}_{j,k}f\right\|_{L^{q}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\tilde{s}_{B,\epsilon}(p,q,q)k}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Thus, for all $1\leq r<p\leq q\leq\infty$,
(4.8) $\left\|\widetilde{T}^{1}_{j,k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $s\geq\tilde{s}_{C,\epsilon}(p,q,r)$ and
$\tilde{s}_{C,\epsilon}(p,q,r):=\min\left\{\tilde{s}_{B,\epsilon}(p,q,p),\tilde{s}_{B,\epsilon}(p,q,q)\right\}.$
Notice that
$\tilde{s}_{C,0}(p,q,r)=\min\left\{\tilde{s}_{A,0}(p,p,p)+2\left(\frac{1}{p}-\frac{1}{q}\right),\tilde{s}_{A,0}(p,q,q)\right\}.$
Consequently, if $(\frac{1}{p},\frac{1}{q})$ satisfies $\tilde{s}_{C,0}(p,q,r)<0$, we establish the desired decay time-space estimate for $\widetilde{T}^{1}_{j,k}$, and the corresponding time-space estimate for $T$ follows.
Remark 4.1.
For any $\epsilon>0$, a computation yields
[…]
If $\tilde{s}_{C,0}(p,q,r)<0$, then (4.8) holds with a decay in $k$, where $(\frac{1}{p},\frac{1}{q})$ belongs to the region with vertices $(0,0)$, $(\frac{1}{4},0)$, $(\frac{3}{8},\frac{1}{8})$, $(\frac{1}{2},\frac{1}{6})$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$. Furthermore, the point $(\frac{3}{8},\frac{1}{8})$ can be removed by interpolation between $(\frac{1}{4},0)$ and $(\frac{1}{2},\frac{1}{6})$. Thus, (4.8) holds with some $s<0$ if $(\frac{1}{p},\frac{1}{q})$ satisfies
$\begin{cases}\frac{1}{q}>\frac{2}{3p}-\frac{1}{6},&\text{for }q\geq 3p^{\prime};\\ \frac{1}{q}>\frac{1}{p}-\frac{1}{3},&\text{for }p^{\prime}<q<3p^{\prime};\\ \frac{1}{q}>\frac{2}{p}-1,&\text{for }q\leq p^{\prime}.\end{cases}$
Combining $1\leq r<p\leq q\leq\infty$ with the necessary condition $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$, the region of $(\frac{1}{p},\frac{1}{q})$ in Theorem 1.4 for this case is illustrated in Figure 5. Here, we assume $1\leq r\leq\frac{3}{2}$, and the figures for other ranges of $r$ can be constructed similarly.
Figure 5: The region of boundedness in (1.4) for the case $1\leq r<p\leq q\leq\infty$ and $1\leq r\leq\frac{3}{2}$.
4.2 Necessary Part
In this subsection, we establish necessary conditions. We begin with a lemma analogous to Lemma 3.2.
Lemma 4.2.
Suppose $G(x,u)\in C_{c}^{1}\left(\frac{1}{2},\frac{5}{2}\right)$ with respect to $u$ for all $x\in\mathbb{R}^{2}$. Then, for any $1\leq q,r\leq\infty$, the following inequality holds:
$\left\|G\right\|_{L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim\left\|G\right\|^{\frac{1}{r^{\prime}}}_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\left\|\partial_{u}G\right\|^{\frac{1}{r}}_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}.$
Proof.
It suffices to consider $r\in[1,\infty)$. Observe that for all $u\in\left(\frac{1}{2},\frac{5}{2}\right)$,
$\left\|G(x,u)\right\|_{L_{x}^{q}(\mathbb{R}^{2})}^{r}\lesssim\int_{\mathbb{R}}\left\|G(x,v)\right\|_{L_{x}^{q}(\mathbb{R}^{2})}^{r-1}\left|\partial_{v}\|G(x,v)\|_{L_{x}^{q}(\mathbb{R}^{2})}\right|\,\mathrm{d}v.$
Applying Hölder’s inequality, we obtain
$\left\|G(x,u)\right\|_{L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}^{r}\lesssim\left\|G(x,u)\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}^{r-1}\left\|\partial_{u}\|G(x,u)\|_{L_{x}^{q}(\mathbb{R}^{2})}\right\|_{L^{r}_{u}(\mathbb{R})}.$
Combining this with the inequality
$\left|\partial_{u}\|G(x,u)\|_{L_{x}^{q}(\mathbb{R}^{2})}\right|\leq\|\partial_{u}G(x,u)\|_{L_{x}^{q}(\mathbb{R}^{2})}$
for all $u\in(\frac{1}{2},\frac{5}{2})$, which follows from Hölder’s inequality, yields the desired estimate. This completes the proof. ∎
Now, assume that
(4.9) $\left\|Tf\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$
holds for some $1\leq p,q,r\leq\infty$.
4.2.1 Proof of (I) in Theorem 1.4
We restrict our attention to the case $q\in[1,\infty)$. Let $f\in C_{c}^{\infty}(\mathbb{R}^{2})$ with $\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}=1$. Since $T$ is translation-invariant for any fixed $u\in[1,2]$, it follows from (4.9) that
$\left\|Tf(x-h,u)+Tf(x,u)\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\leq\Theta\|f(x-h)+f(x)\|_{L_{x}^{p}(\mathbb{R}^{2})},$
where $h\in\mathbb{R}^{2}$ and $\Theta:=\sup_{f\in C_{c}^{\infty}(\mathbb{R}^{2}),\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}=1}\|Tf\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}<\infty$. Letting $|h|\to\infty$ and applying Grafakos [13, Exercise 2.5.1], we deduce
$2^{\frac{1}{q}}\left\|Tf(x,u)\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\leq\liminf_{|h|\to\infty}\left\|Tf(x-h,u)+Tf(x,u)\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\leq 2^{\frac{1}{p}}\Theta\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Thus, $\Theta\leq 2^{\frac{1}{p}-\frac{1}{q}}\Theta$, which is impossible unless $q\geq p$ or $T$ is the zero operator. This contradiction establishes $\frac{1}{q}\leq\frac{1}{p}$, completing the proof of (I) of Theorem 1.4.
4.2.2 Proof of (II) of Theorem 1.4
As in Subsection 3.2, if (4.9) holds, then
$\left\|Tf\right\|_{L^{1}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Furthermore,
$\int_{1}^{2}\left|\frac{S_{t_{i}}}{2}\right|\,\mathrm{d}u\leq\left\|\frac{1}{\delta t_{i}}\int_{\frac{S_{t_{i}}}{2}}T\chi_{S_{t_{i}}}(x,u)\,\mathrm{d}x\right\|_{L^{1}_{u}([1,2])}\lesssim\frac{|S_{t_{i}}|^{\frac{1}{q^{\prime}}}}{t_{i}}\left\|T\chi_{S_{t_{i}}}\right\|_{L^{1}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\frac{|S_{t_{i}}|^{\frac{1}{q^{\prime}}}}{t_{i}}\left\|\chi_{S_{t_{i}}}\right\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
Using the facts that $|S_{t_{i}}|\approx t_{i}|\gamma(t_{i})|$ and $\|\chi_{S_{t_{i}}}\|_{L_{x}^{p}(\mathbb{R}^{2})}\approx(t_{i}|\gamma(t_{i})|)^{\frac{1}{p}}$, we obtain (3.24). This implies $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)\geq 0$, proving (II) of Theorem 1.4.
4.2.3 Proof of (III) of Theorem 1.4
This proof is inspired by examples in [31]. For any $u\in[1,2]$ and $\epsilon>0$, let $D_{u}$ be the $\epsilon$-neighborhood of the curve $(ut,u\gamma(t))$ with $t\in(0,1]$. As in [31], there exists $\delta\in(0,\infty)$ such that $T(\chi_{B(0,\delta\epsilon)})(x,u)\geq\epsilon\chi_{D_{u}}(x)$. Since $|D_{u}|\approx\epsilon$ for all $u\in[1,2]$, we have
$\left\|T(\chi_{B(0,\delta\epsilon)})\right\|_{L_{x}^{q}(\mathbb{R}^{2})}\geq\epsilon\left\|\chi_{D_{u}}\right\|_{L_{x}^{q}(\mathbb{R}^{2})}\gtrsim\epsilon^{1+\frac{1}{q}},$
which implies
$\left\|T(\chi_{B(0,\delta\epsilon)})\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\gtrsim\epsilon^{1+\frac{1}{q}}.$
Combining this with $\|Tf\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$ and $\|\chi_{B(0,\delta\epsilon)}\|_{L_{x}^{p}(\mathbb{R}^{2})}\approx\epsilon^{\frac{2}{p}}$, we obtain $\epsilon^{\frac{1}{q}-\frac{2}{p}+1}\lesssim 1$ for all $\epsilon>0$. For sufficiently small $\epsilon$, this requires $\frac{1}{q}\geq\frac{2}{p}-1$.
On the other hand, for $\epsilon>0$, define
$Q:=\left\{x\in\mathbb{R}^{2}:|x\cdot e_{1}|\leq 6\epsilon\text{ and }|x\cdot e_{2}|\leq[3+2|\gamma^{\prime\prime}(1)|]\epsilon^{2}\right\},$
where $e_{1}:=\frac{(-1,-\gamma^{\prime}(1))}{\sqrt{1+\gamma^{\prime}(1)^{2}}}$ and $e_{2}:=\frac{(-\gamma^{\prime}(1),1)}{\sqrt{1+\gamma^{\prime}(1)^{2}}}$. Additionally, define
$\Omega_{u_{i}}:=\left\{x\in\mathbb{R}^{2}:\left|\left(x-(u_{i},u_{i})\right)\cdot e_{1}\right|\leq\epsilon\text{ and }\left|\left(x-(u_{i},u_{i})\right)\cdot e_{2}\right|\leq\epsilon^{2}\right\},$
where $\{u_{i}\}\subset[1,2]$ satisfies $|u_{i+1}-u_{i}|=\frac{4\epsilon^{2}}{|(1,1)\cdot e_{2}|}$ for all $i$. Moreover, for sufficiently small $\epsilon>0$, we obtained in [31] that
$T\chi_{Q}(x,u)\geq\int_{1-\frac{\epsilon}{\sqrt{1+\gamma^{\prime}(1)^{2}}}}^{1}\chi_{Q}(x_{1}-ut,x_{2}-u\gamma(t))\,\mathrm{d}t=\frac{\epsilon}{\sqrt{1+\gamma^{\prime}(1)^{2}}}$
for all $x\in\Omega_{u_{i}}$ and $u\in(u_{i}-\epsilon^{2},u_{i}+\epsilon^{2})$. Since $\Omega_{u_{i}}\cap\Omega_{u_{i^{\prime}}}=\emptyset$ for $i\neq i^{\prime}$ and the number of elements in $\{u_{i}\}\subset[1,2]$ is equivalent to $\epsilon^{-2}$, we conclude
$\left\|T\chi_{Q}\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\geq\left\{\sum_{i}\int_{u_{i}-\epsilon^{2}}^{u_{i}+\epsilon^{2}}\left[\int_{\Omega_{u_{i}}}\left|T\chi_{Q}\right|^{q}\,\mathrm{d}x\right]^{\frac{r}{q}}\,\mathrm{d}u\right\}^{\frac{1}{r}}\gtrsim\epsilon^{1+\frac{3}{q}}.$
Combining this with $\|\chi_{Q}\|_{L^{p}(\mathbb{R}^{2})}\approx\epsilon^{\frac{3}{p}}$ and $\|Tf\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times[1,2])}\lesssim\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}$, we obtain $\epsilon^{1+\frac{3}{q}-\frac{3}{p}}\lesssim 1$ for all sufficiently small $\epsilon>0$. Letting $\epsilon\to 0$ implies $\frac{1}{q}\geq\frac{1}{p}-\frac{1}{3}$ for all $r\in[1,\infty]$.
Thus, we obtain $\frac{1}{q}\geq\max\{\frac{2}{p}-1,\frac{1}{p}-\frac{1}{3}\}$, and (III) of Theorem 1.4 is proved.
4.2.4 Proof of (IV) of Theorem 1.4
From (4.3), setting $\bar{s}_{2,-\frac{1}{2}}(p,q)<0$ yields a decay estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$, where $(\frac{1}{p},\frac{1}{q})$ lies in the open region with vertices $(0,0)$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$. Then, $T:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ holds on the intersection of this region with $1+(1+\omega)\left(\frac{1}{q}-\frac{1}{p}\right)>0$. Thus, $T:L_{x}^{p}(\mathbb{R}^{2})\to L_{x}^{q}(\mathbb{R}^{2})$ holds on this region uniformly in $u\in[1,2]$. Combining this with [31, Theorem 2.1], we conclude that $(0,0)$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$ are the endpoints of the estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$. Therefore, $\bar{s}_{2,-\frac{1}{2}}(p,q)$ is the infimum ensuring (4.3) holds.
On the other hand, as in Subsection 5.1, for $1\leq p\leq q\leq\infty$, the $L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ estimates for $E_{j,k}-F_{j,k}$ and $\widetilde{T}^{1,a}_{j,k}$ are bounded by $2^{-Nk}$ for large $N$. Recall that $Q_{k}f=(\varrho_{k}(\xi)\hat{f}(\xi))^{\vee}$, and a similar estimate to (4.3) holds. Namely, if
$\left\|\widetilde{T}^{1}_{j,k}Q_{k}f\right\|_{L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
then $s\geq\bar{s}_{2,-\frac{1}{2}}(p,q)$. Now, from Lemma 4.2 and (4.9), we have
$\left\|\widetilde{T}^{1}_{j,k}Q_{k}f\right\|_{L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\frac{1}{r}k}\left\|\widetilde{T}^{1}_{j,k}Q_{k}f\right\|_{L^{r}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\frac{1}{r}k}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
which implies $\frac{1}{r}\geq\bar{s}_{2,-\frac{1}{2}}(p,q)$, i.e.,
$\frac{1}{q}\geq\max\left\{\frac{1}{2p}-\frac{1}{2r},\frac{2}{p}-\frac{r+1}{r}\right\}.$
From (4.2), setting $s_{-\frac{1}{2},0}(p,q,q)<0$ yields the decay estimate $F_{j,k}:L_{x}^{p}(\mathbb{R}^{2})\to L^{\infty}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})$ when $(\frac{1}{p},\frac{1}{q})$ lies in the open region with vertices $(0,0)$, $(\frac{1}{2},\frac{1}{6})$, $(\frac{2}{3},\frac{1}{3})$, and $(1,1)$. This result is nearly sharp by [31]. Similarly, if
$\left\|\widetilde{T}^{1}_{j,k}Q_{k}f\right\|_{L^{q}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{sk}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})},$
then $s\geq s_{-\frac{1}{2},0}(p,q,q)$. On the other hand, by (4.9), applying the Sobolev embedding theorem with respect to $u$ for $q>r$ and $\textrm{H}\ddot{\textrm{o}}\textrm{lder}$’s inequality for $q\leq r$ yields
$\left\|\widetilde{T}^{1}_{j,k}Q_{k}f\right\|_{L^{q}_{u}L_{x}^{q}(\mathbb{R}^{2}\times\mathbb{R})}\lesssim 2^{\max\left\{\frac{1}{r}-\frac{1}{q},0\right\}k}\|f\|_{L_{x}^{p}(\mathbb{R}^{2})}.$
It further yields $\max\{\frac{1}{r}-\frac{1}{q},0\}\geq s_{-\frac{1}{2},0}(p,q,q)$. Note that $s_{-\frac{1}{2},0}(p,q,q)=\tilde{s}_{A,0}(p,q,q)$, we obtain
$\max\left\{\frac{1}{r}-\frac{1}{q},0\right\}\geq\tilde{s}_{A,0}(p,q,q).$
Putting these results together, (IV) of Theorem 1.4 is proved.
Acknowledgments. The authors would like to thank N. Liu for helpful suggestions and discussions. J. Li is supported by National Natural Science Foundation of China (No. 12471090). Z. Lou is supported by National Natural Science Foundation of China (No. 12471093), Guangdong Basic and Applied Basic Research Foundation (No. 2024A1515010468), and LKSF STU-GTIIT Joint-research Grant (No. 2024LKSF06). H. Yu is supported by National Natural Science Foundation of China (No. 12201378) and Guangdong Basic and Applied Basic Research Foundation (No. 2023A1515010635).